### v1

Abstract
We study the application of a quasi-Monte Carlo (QMC) method to a class of semi-linear parabolic reaction-diffusion partial differential equations used to model tumor growth. Mathematical models of tumor growth are largely phenomenological in nature, capturing infiltration of the tumor into surrounding healthy tissue, proliferation of the existing tumor, and patient response to therapies, such as chemotherapy and radiotherapy. Considerable inter-patient variability, inherent heterogeneity of the disease, sparse and noisy data collection, and model inadequacy all contribute to significant uncertainty in the model parameters. It is crucial that these uncertainties can be efficiently propagated through the model to compute quantities of interest (QoIs), which in turn may be used to inform clinical decisions. We show that QMC methods can be successful in computing expectations of meaningful QoIs. Well-posedness results are developed for the model and used to show a theoretical error bound for the case of uniform random fields. The theoretical linear error rate, which is superior to that of standard Monte Carlo, is verified numerically. Encouraging computational results are also provided for lognormal random fields, prompting further theoretical development.
Key words. uncertainty quantification, quasi-Monte Carlo methods, semi-linear parabolic equations, nonlinear analysis, computational oncology
MSC codes. 65D30, 65D32, 92B05, 92C50, 35K58
1 Introduction
Uncertainty quantification (UQ) is crucial in precision medicine, establishing trust in computational models for high-consequence applications. This is particularly salient in oncology, where biomedical imaging plays a critical role in the diagnosis and management of a variety of tumors. However, these data are noisy, infrequently collected, and only indirectly informative, all contributing to significant uncertainty in calibrated models. Practical decision making requires efficient methods for the forward propagation of this uncertainty through mathematical and computational models as well as the estimation of clinically relevant quantity of interests (QoIs). In this work, we develop quasi-Monte Carlo (QMC) methods for efficiently estimating the expectation of QoIs from non-linear partial differential equation (PDE) models of tumor growth.
We consider a semi-linear parabolic reaction-diffusion PDE model of tumor growth that captures both invasion into surrounding tissue [1] and proliferation of the existing lesion [5]. This is an established model in the field of computational oncology [44] and has been extended to account for patient response to systemic chemo-radiation therapy [25]. Here, we model the uncertain diffusion and proliferation coefficients as random fields, the uncertainty in which leads to uncertainty in the modeled state and QoIs. We will consider the expectation to be integrated over finitely many random variables by truncating a series expansion representing the formally infinite-dimensional input random fields to a sum over finitely many terms. To efficiently address the high-dimensional integral, we will employ QMC methods with the aim of improving upon the slow convergence of standard Monte Carlo (MC) simulation.
QMC methods are a class of quadrature rules designed to compute high-dimensional integrals (or equivalently expectations). The key benefits of QMC methods are that they can be tailored to achieve dimension-independent error bounds with faster than MC convergence rates for smooth integrands. QMC methods have demonstrated impressive performance in several UQ applications involving PDEs with random coefficients or domains, such as linear elliptic PDEs [12, 15, 18, 19, 20, 22, 29], elliptic eigenvalue problems [17, 37] and more recently linear parabolic equations with control [21]. However, much of the theory has focused on linear PDEs. The present work extends the theory in the case of non-linear, specifically semi-linear, parabolic equations, providing a roadmap for adapting QMC methods to wider classes of applications. Apart from the above work on eigenvalue problems [^3] While eigenvalue problems are also nonlinear, the differential operators [17, 37] are linear., to the best of our knowledge this paper is the first to analyze QMC methods for a non-linear PDE.
Well-posedness results are developed for the parametric PDE model in general. Rigorous error estimates are proven for the special case of uniform random fields and confirmed with numerical experiments. Note that we do not analyze the error originating from the spatio-temporal discretization of the governing PDE, leaving this for future work. Encouraging numerical results are also presented for a more realistic test case of log-normal random fields on a complex, unstructured domain, used previously in models of gliomas in the brain [30, 38].
One of our key intermediate results in the QMC error analysis is a bound on the derivatives of the solution with respect to uniform random parameters. Similar parametric regularity results have also been studied for related classes of semi-linear elliptic PDEs in [7, 23, 24]; see also e.g., [2, 8] for other nonlinear PDE or SPDE.
However, the nonlinear term considered here, namely, u(1u)u(1-u) (see (2.1)), does not fit the settings in [7, 23, 24]. In particular, it is not monotone and so requires a novel proof, which uses the property that if our initial condition lies in $[0,1]$ then the solution remains in $[0,1]$ (proven in Section 2.5). We also makes use of the falling factorial combinatorial bound from [6]. We stress that obtaining regularity bounds for parabolic (time-dependent) nonlinear problems here is significantly harder than for elliptic problems in [6, 7, 23, 24] due to the need to handle both space and time derivatives and integrals in the norms. The parametric regularity bounds presented here are also of independent interest beyond QMC since they are often required to analyze other parametric approximation methods, e.g., sparse grids or stochastic collocation.
The remainder of this paper is organized as follows. In Section 2 we present the governing equations and variational framework, and then establish well-posedness and an a priori bound with explicit constant depending on the parameters for general random field models. A parametric regularity bound is obtained for the case of uniform random fields in Section 3. Our main result addresses the QMC error rate in Section 4. The theoretical results are confirmed with supporting numerical experiments in Section 5. Conclusions are drawn in Section 6.
2 Problem formulation
2.1 Governing equation
The invasion of a solid tumor into surrounding healthy tissue is typically modeled as a diffusion process, while the proliferation of existing tumor is modeled with logistic growth subject to some carrying capacity. These phenomenological assumptions give rise to a semi-linear parabolic reaction-diffusion PDE [36], which has demonstrated predictive value for a variety of solid tumors [9, 26, 27, 33]. The model is given by
$\mbox{PDE I:}\quad\begin{cases}\displaystyle\frac{\partial u}{\partial t}-\nabla\cdot(a\,\nabla u)-\kappa\,u\,(1-u)+f\,u=0,\qquad&{\boldsymbol{x}}\in\Omega,\quad t\in I,\\ u({\boldsymbol{x}},0)=u_{0}({\boldsymbol{x}}),&{\boldsymbol{x}}\in\Omega\\ \nabla u\cdot{\boldsymbol{n}}=0,&{\boldsymbol{x}}\in\partial\Omega,\quad t\in I,\end{cases}$ (2.1)
where $u=u^{\boldsymbol{y}}({\boldsymbol{x}},t)$ models the tumor volume fraction varying in time $t$ on the interval $I:=[0,T]$ and over spatial coordinates ${\boldsymbol{x}}\in\Omega$, where $\Omega\subset\mathbb{R}^{d}$ is the bounded spatial domain with physical dimension $d\in\{2,3\}$ and Lipschitz boundary $\partial\Omega$ with outward unit normal $\boldsymbol{n}$}. The homogeneous Neumann boundary condition ensures that the tumor does not grow beyond the boundary of the domain. The linear reaction term fuf\,u models the effect of treatment. The initial tumor volume fraction $u_{0}({\boldsymbol{x}})$ may be estimated from medical imaging data. It is natural to assume
$0\leq u_{0}({\boldsymbol{x}})\leq 1\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega.$ (2.2)
Systemic therapies, such as chemotherapy and radiotherapy form the basis of many treatment protocols, for example, gliomas in the brain [43]. Combination chemoradiation may be modeled as a linear effect on the state, $f=f({\boldsymbol{x}},t)=f_{\rm rt}({\boldsymbol{x}},t)+f_{\rm ct}({\boldsymbol{x}},t)$, with
$f_{\rm rt}({\boldsymbol{x}},t)$ $=\gamma_{\text{rt},\varepsilon}\sum_{k}\big[1-\exp(-\alpha_{\text{rt}}\,z_{\text{rt},k}({\boldsymbol{x}})-\beta_{\text{rt}}\,z^{2}_{\text{rt},k}({\boldsymbol{x}}))\big]\,\,\mathbbold{1}(t\in[t_{\text{rt},k},t_{\text{rt},k}+\varepsilon]),$ (2.3)
$f_{\rm ct}({\boldsymbol{x}},t)$ $=\alpha_{\rm ct}\,\sum_{\ell}z_{\text{ct},\ell}({\boldsymbol{x}})\,\exp(-\beta_{\rm ct}(t-\tau_{{\rm ct},\ell}))\,\mathbbold{1}(t\geq\tau_{\text{ct},\ell}),$ (2.4)
where ${\mathcal{T}}_{\rm rt}=\{\tau_{{\rm rt},k}\}_{k}$ and ${\mathcal{T}}_{\rm ct}=\{\tau_{{\rm ct},\ell}\}_{\ell}$ denote the collection of times at which radiotherapy and chemotherapy are administered, respectively, and $\mathbbold{1}$ denotes the indicator function. Radiotherapy is assumed to be instantaneous at the moment of treatment with no lasting or time-delaying effects. The application duration is much smaller than the time-discretization, that is, $\varepsilon\ll\Delta t$, where $\Delta t$ is the step size. The surviving fraction is computed with the linear-quadratic model: $z_{{\rm rt},k}({\boldsymbol{x}})$ is the radiotherapy dosage, $\alpha_{\rm rt}>0$ and $\beta_{\rm rt}>0$ are parameters describing the radiosensitivity of the tissue, and $\gamma_{{\rm rt},\varepsilon}>0$ is a parameter that ensures appropriate physical units [3, 35, 40]. Chemotherapy is assumed to follow a decaying exponential model: $z_{{\rm ct},\ell}({\boldsymbol{x}})$ is the drug concentration, $\alpha_{\rm ct}>0$ is the efficacy, and $\beta_{\rm ct}>0$ is the clearance rate [25, 27].
The superscript $\boldsymbol{y}$} denotes a parametric vector modeling any randomness and we allow time-varying coefficients for the diffusion $a=a^{\boldsymbol{y}}({\boldsymbol{x}},t)$ and proliferation $\kappa=\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)$. The precise details on the modeling of the random fields is deferred to Section 3. However, we will restrict the parametric vectors $\boldsymbol{y}$} to some admissible parameter set $U\subset{\mathbb{R}}^{\mathbb{N}}$, such that
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}<\infty,\\ 0=f_{\min}\leq f({\boldsymbol{x}},t)\leq f_{\max}<\infty,\end{cases}\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.5)
2.2 Functional framework
Let $H^{1}(\Omega)$ denote the usual first-order Sobolev space on $\Omega$ with (squared) norm
$\|v\|_{H^{1}(\Omega)}^{2}:=\|v\|_{L^{2}(\Omega)}^{2}+\|\nabla v\|_{L^{2}(\Omega)}^{2}.$
Let $H^{1}(\Omega)^{*}$ denote the dual space of $H^{1}(\Omega)$, i.e., the space of all bounded linear functionals on $H^{1}(\Omega)$. We denote the duality pairing between $H^{1}(\Omega)^{*}$ and $H^{1}(\Omega)$ by $\langle\cdot,\cdot\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}$, with $L^{2}(\Omega)=L^{2}(\Omega)^{*}$ as the pivot space. The dual norm is defined by
$\|z\|_{H^{1}(\Omega)^{*}}:=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle z,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}z\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}.$
It is advantageous to consider functions of time with values in Banach spaces when dealing with time-dependent problems. For a Banach space $B$, let the space $L^{2}(I;B)$ consist of all measurable functions $v:I\to B$ with finite norm $\|v\|_{L^{2}(I;B)}^{2}:=\int_{I}\|v(\cdot,t)\|_{B}^{2}\,{\mathrm{d}}t$, and let $C(I;B)$ consist of all functions $v:I\to B$ that are continuous at every $t\in I$ with norm $\|v\|_{C(I;B)}:=\max_{t\in I}\|v(\cdot,t)\|_{B}$. Specifically, in this paper we make use of the following norms:
$\|v\|_{L^{2}(I;L^{2}(\Omega))}^{2}$ $:=\int_{I}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t,$
$\|v\|_{L^{2}(I;H^{1}(\Omega))}^{2}$ $:=\int_{I}\|v(\cdot,t)\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t=\int_{I}\big(\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\nabla v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\big)\,{\mathrm{d}}t,$
$\|v\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}$ $:=\int_{I}\|v(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\,{\mathrm{d}}t,$
$\|v\|_{C(I;L^{2}(\Omega))}$ $:=\max_{t\in I}\|v(\cdot,t)\|_{L^{2}(\Omega)}.$
Note that $L^{2}(I;H^{1}(\Omega))^{*}=L^{2}(I;H^{1}(\Omega)^{*})$.
We follow Tröltzsch [45, §3.4] to analyze our PDE solution in the space $\mathcal{X}$} (denoted by $W(0,T)$ in [45]) which consists of all functions $v\in L^{2}(I;H^{1}(\Omega))$ with (distributional) derivative $\tfrac{\partial}{\partial t}v\in L^{2}(I;H^{1}(\Omega)^{*})$, and we equip the space $\mathcal{X}$} with the (graph) norm
$\|v\|_{\mathcal{X}}^{2}$ $:=\|v\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}v\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}$
$\,=\int_{I}\big(\|v(\cdot,t)\|_{H^{1}(\Omega)}^{2}+\|\tfrac{\partial}{\partial t}v(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t$
$\,=\int_{I}\big(\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\nabla v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\tfrac{\partial}{\partial t}v(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t.$
A fundamental property we will use is that the space $\mathcal{X}$} is embedded in $C(I;L^{2}(\Omega))$, see [45, Theorems 3.10 and 3.11], and therefore we formally write for $v\in{\mathcal{X}}$,
$\int_{I}\int_{\Omega}\frac{\partial v}{\partial t}v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|v(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
We will make use of a number of embedding constants which we summarize together here:
$\|v\|_{H^{1}(\Omega)^{*}}$ $\leq\|v\|_{L^{2}(\Omega)}\leq\|v\|_{H^{1}(\Omega)},$
$\|v\|_{L^{4}(\Omega)}$ $\leq\theta_{1}\,\|v\|_{H^{1}(\Omega)},$ (2.6)
$\|v\|_{C(I;L^{2}(\Omega))}$ $\leq\theta_{2}\,\|v\|_{\mathcal{X}}.$ (2.7)
The embedding (2.6) is due to the Sobolev embedding theorem for $H^{1}(\Omega)$, see e.g., [45, Theorem 7.1], which state that $H^{1}(\Omega)$ is embedded in $L^{p}(\Omega)$ when $d=2$ and $1\leq p<\infty$, and when $d\geq 3$ and $1\leq p\leq\frac{2d}{d-2}$.
Another property we will use is that the space $H^{1}(\Omega)$ is closed under addition, subtraction, and the maximum and minimum operations, see e.g., [16, Lemma 7.6]. This allows us to define continuous piecewise-smooth test functions in $H^{1}(\Omega)$.
2.3 Reparameterization
In the case where $f=0$, that is, untreated growth, the PDE model (2.1) is the well-studied Fisher-KPP equation [14, 28], and well-posedness results are readily available, see e.g., [39, §3.10.3]. However, we consider $f\neq 0$, thus standard results are not immediately applicable. PDE I (2.1) can be rewritten as
$\frac{\partial u}{\partial t}-\nabla\cdot(a\,\nabla u)+\kappa\,u^{2}+(f-\kappa)\,u=0.$
It is not known a priori whether $f\geq\kappa$ holds. Moreover, we are unable to deduce monotonicity of the function $\kappa\,u^{2}+(f-\kappa)\,u$ with respect to $u$. Consequently, we cannot claim the existence and uniqueness of a solution using known results, such as those in [45].
Following a standard technique for treating parabolic PDEs involving a linear term with a negative coefficient (see e.g., [13, §7.5 Problem 8]), we use the substitution
$u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w^{\boldsymbol{y}}({\boldsymbol{x}},t)\,$,
for some $\lambda^{\boldsymbol{y}}\in{\mathbb{R}}$, possibly depending on $\boldsymbol{y}$}, to obtain
$u=e^{\lambda\,t}\,w\quad\implies\quad\frac{\partial u}{\partial t}=\lambda\,e^{\lambda\,t}\,w+e^{\lambda\,t}\,\frac{\partial w}{\partial t}\quad\mbox{and}\quad\nabla u=e^{\lambda\,t}\,\nabla w,$
where to simplify notation we have omitted the $\boldsymbol{y}$} dependence. Therefore, after dividing through by $e^{\lambda\,t}$ and regrouping, the PDE becomes
$\frac{\partial w}{\partial t}-\nabla\cdot(a\,\nabla w)+\kappa\,e^{\lambda\,t}\,w^{2}+(\lambda+f-\kappa)\,w=0.$
Thus we have the reparametrized PDE problem,
$\mbox{PDE II:}\quad\begin{cases}\displaystyle\frac{\partial w}{\partial t}-\nabla\cdot(a\,\nabla w)+b\,w^{2}+c\,w=0,\qquad&{\boldsymbol{x}}\in\Omega,\quad t\in I,\\ w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}}),&{\boldsymbol{x}}\in\Omega\\ \nabla w\cdot{\boldsymbol{n}}=0,&{\boldsymbol{x}}\in\partial\Omega,\quad t\in I,\end{cases}$ (2.8)
with coefficients and initial condition related to PDE I (2.1) via
$\begin{cases}b^{\boldsymbol{y}}({\boldsymbol{x}},t):=\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\,\exp(\lambda^{\boldsymbol{y}}\,t),\\ c^{\boldsymbol{y}}({\boldsymbol{x}},t):=\lambda^{\boldsymbol{y}}+f({\boldsymbol{x}},t)-\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t),\\ w_{0}({\boldsymbol{x}}):=u_{0}({\boldsymbol{x}}).\end{cases}$ (2.9)
To ensure that $c^{\boldsymbol{y}}({\boldsymbol{x}},t)>0$, we may take any $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}$. From (2.2) we have $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$. It follows from (2.5) that
$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq b^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}\,e^{\lambda^{\boldsymbol{y}}\,T}<\infty,\\ 0<\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}}\leq c^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\lambda^{\boldsymbol{y}}+f_{\max}-\kappa_{\min}^{\boldsymbol{y}}<\infty,\end{cases}\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.10)
2.4 Well-posedness
In this section, we provide an existence and uniqueness result for the system. We begin by deriving the parametric variational formulation of the PDE problem.
For fixed ${\boldsymbol{y}}\in U$, a function $u^{\boldsymbol{y}}\in{\mathcal{X}}$ is said to be a weak solution for PDE I (2.1) if
$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial u}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla u\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(\kappa\,u^{2}+(f-\kappa)\,u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}u(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}u_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.11)
The first condition was obtained by multiplying the PDE with a test function $v\in{\mathcal{X}}$, integrating over ${\boldsymbol{x}}\in\Omega$ and $t\in I$, applying integration by parts on the diffusion term, and noting that $\int_{\partial\Omega}a\,\nabla u\cdot{\boldsymbol{n}}\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ due to the Neumann boundary condition.
Analogously, a function $w^{\boldsymbol{y}}\in{\mathcal{X}}$ is said to be a weak solution for PDE II (2.8) if
$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}w(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}w_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.12)
Theorem 2.1 (existence, uniqueness, boundedness)
For every ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, consider PDE II (2.8) with bounded initial condition $w_{0}\in L^{2}(\Omega)$, $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.9)–(2.10). Then, there exists a unique weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE II, which is bounded by $0\leq w^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$, and it satisfies the a priori bound $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
In turn, $u^{\boldsymbol{y}}=e^{\lambda^{\boldsymbol{y}}\,t}\,w^{\boldsymbol{y}}\in{\mathcal{X}}$ is the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}=w_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5), and it is bounded by $0\leq u^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq 1$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
Before proving this theorem, we first remark that there is considerable theory developed for the case of monotone semi-linear equations [45], which we could readily make use of were the term $b\,w^{2}+c\,w$ of PDE II (2.8) monotone in $w$. Unfortunately, $b\,w^{2}+c\,w$ is generally not monotone, unless it is known that $w\geq 0$. Moreover, recall that the solution for PDE I (2.1) represents the tumor volume fraction and should naturally satisfy $0\leq u\leq 1$, corresponding to $0\leq w\leq e^{-\lambda\,t}$. But these bounds are not immediately obvious. We therefore modify our PDE II by imposing a restriction on the term $b\,w^{2}+c\,w$ whenever $w<0$ or $w>e^{-\lambda\,t}$:
[…]
where we define a continuous piecewise cutoff function
$d(w):=\begin{cases}0&\mbox{if }w<0,\\ b\,w^{2}+c\,w&\mbox{if }0\leq w\leq e^{-\lambda\,t},\\ b\,e^{-2\lambda\,t}+c\,e^{-\lambda\,t}=(\lambda+f)\,e^{-\lambda\,t}&\mbox{if }w>e^{-\lambda\,t}.\end{cases}$ (2.14)
A weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE III (2.13) satisfies (2.12) with $b\,w^{2}+c\,w$ replaced by $d(w)$.
Our argument for proving Theorem 2.1 goes as follows:
  1. 1.
    For fixed ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}$, we infer from [45, Assumptions 5.1 and 5.2, Lemmas 5.3 and 7.10] that there exists a unique weak solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III with bounded coefficients (2.10), with a priori bound $\|w_{*}^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ for $C^{\boldsymbol{y}}>0$.
  2. 2.
    We prove that if the initial condition satisfies $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$ then this unique solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III satisfies $0\leq w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
  3. 3.
    Then, for this unique solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$, PDE III holds with $d(w)$ defined only by the middle case. This shows that $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ is also a solution for PDE II. This confirms the existence of a solution for PDE II satisfying $0\leq w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$, given the bounded initial condition $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$.
  4. 4.
    Finally we prove that this solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ for PDE II is unique, and hence conclude that $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)$ is the unique solution of PDE I, and it satisfies $0\leq u({\boldsymbol{x}},t)\leq 1$ for a.e. ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
The strategy of working with a cutoff function in PDE III and using piecewise test functions (see our proof below) appears already in the literature, sometimes referred to as Stampacchia truncation, see e.g., [4, 10, 39, 45]. However, the problem settings in those references differ from ours, so we include a proof for completeness.
[Proof of Theorem 2.1] We begin by proving the existence of a solution.
Existence. We fix ${\boldsymbol{y}}\in U$ and assume bounded coefficients (2.10). We write the cutoff function (2.14) showing explicit dependence on all arguments,
$d(w)=d({\boldsymbol{x}},t,w):=\begin{cases}0&\mbox{if }w^{\boldsymbol{y}}(x,t)<0,\\ b^{\boldsymbol{y}}({\boldsymbol{x}},t)\,\big(w^{\boldsymbol{y}}({\boldsymbol{x}},t)\big)^{2}+c^{\boldsymbol{y}}({\boldsymbol{x}},t)\,w^{\boldsymbol{y}}(x,t)&\mbox{if }0\leq w^{\boldsymbol{y}}(x,t)\leq e^{-\lambda^{\boldsymbol{y}}\,t},\\ b^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{-2\lambda^{\boldsymbol{y}}\,t}+c^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{-\lambda^{\boldsymbol{y}}\,t}&\mbox{if }w^{\boldsymbol{y}}(x,t)>e^{-\lambda^{\boldsymbol{y}}\,t}.\end{cases}$
We see that PDE III satisfies [45, Assumption 5.1]: $\Omega$ is a bounded Lipschitz domain; for fixed $w\in{\mathbb{R}}$, the function $d({\boldsymbol{x}},t,w)$ is measurable with respect to ${\boldsymbol{x}}\in\Omega$ and $t\in I$; for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is monotonically increasing (nondecreasing) with respect to $w\in{\mathbb{R}}$. We see that PDE III also satisfies [45, Assumption 5.2]: for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is uniformly bounded and Lipschitz in $w\in{\mathbb{R}}$.
Hence, by [45, Lemma 5.3] and subsequent discussions, there exists a unique weak solution of PDE III in $\mathcal{X}$}. Consequently, by [45, Lemma 7.10], this solution satisfies $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
We proceed to prove that this unique weak solution for PDE III is bounded between 0 and $e^{-\lambda\,t}$ if the initial condition is bounded by $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$. Integrating (2.13), using integration by parts on the diffusion term and using the zero Neumann boundary condition, we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}d(w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.15)
Lower bound. We define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ to be the “unwanted lower part” of $w$,
$w_{\#}({\boldsymbol{x}},t):=\min(w({\boldsymbol{x}},t),0)=\begin{cases}w({\boldsymbol{x}},t)&\mbox{if }w({\boldsymbol{x}},t)<0,\\ 0&\mbox{otherwise}.\end{cases}$
For a.e. ${\boldsymbol{x}}\in\Omega$, since $w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}})\geq 0$, we have $w_{\#}({\boldsymbol{x}},0)=0$ and $\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$. If $w({\boldsymbol{x}},t)<0$ then
$\frac{\partial w}{\partial t}\,w_{\#}=\frac{\partial w_{\#}}{\partial t}\,w_{\#},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=0,$ (2.16)
where we used Case 1 of (2.14). On the other hand, if $w({\boldsymbol{x}},t)\geq 0$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.16) hold trivially. Putting (2.16) into (2.15) gives for a.e. $t\in I$,
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}$ is a nonincreasing function of $t$. Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},t)\geq 0\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
Upper bound. Now we define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ on the “unwanted upper part” of $w$,
$w_{\#}({\boldsymbol{x}},t):=\max(w({\boldsymbol{x}},t)-e^{-\lambda\,t},0)=\begin{cases}w({\boldsymbol{x}},t)-e^{-\lambda\,t}&\mbox{if }w({\boldsymbol{x}},t)>e^{-\lambda\,t},\\ 0&\mbox{otherwise}.\end{cases}$
For a.e. ${\boldsymbol{x}}\in\Omega$, since $w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}})\leq 1$, we have $w_{\#}({\boldsymbol{x}},0)=0$ and $\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$. If $w({\boldsymbol{x}},t)>e^{-\lambda\,t}$ then $\frac{\partial w_{\#}}{\partial t}=\frac{\partial w}{\partial t}+\lambda\,e^{-\lambda\,t}$, and so
$\frac{\partial w}{\partial t}\,w_{\#}=\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#},\;\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\;d(w)\,w_{\#}=(\lambda+f)\,e^{-\lambda\,t}\,w_{\#},$ (2.17)
where we used Case 3 of (2.14). On the other hand, if $w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.17) hold trivially. Putting (2.17) into (2.15) gives for a.e. $t\in I$,
$\int_{\Omega}\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(\lambda+f)\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}},$
which yields
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}f\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}$ is a nonincreasing function of $t$. Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
$w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
Uniqueness. We have shown that the unique solution of PDE III satisfies $0\leq w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$ when the initial condition satisfies $0\leq w_{0}({\boldsymbol{x}})\leq 1$. For this solution, PDE III holds with $d(w)$ defined only by Case 2 in (2.14). Thus, this solution for PDE III is also a solution for PDE II, demonstrating the existence of a solution for PDE II satisfying $0\leq w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$.
Suppose that $w_{1}$ and $w_{2}$ are two solutions for PDE II, both bounded between 0 and $e^{-\lambda\,t}$. Then from (2.8) we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w_{1}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{1}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{1}^{2}+c\,w_{1})\,v\,{\mathrm{d}}{\boldsymbol{x}},$
$\int_{\Omega}\frac{\partial w_{2}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{2}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{2}^{2}+c\,w_{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
Subtracting the two equations, writing $w_{1}^{2}-w_{2}^{2}=(w_{1}+w_{2})\,(w_{1}-w_{2})$, and taking $v=w_{1}(\cdot,t)-w_{2}(\cdot,t)\in H^{1}(\Omega)$, we obtain
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}^{2}$
$=-\int_{\Omega}a\,|\nabla(w_{1}-w_{2})|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,(w_{1}+w_{2})+c)\,(w_{1}-w_{2})^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0,$
where we used $w_{1}({\boldsymbol{x}},t)\geq 0$ and $w_{2}({\boldsymbol{x}},t)\geq 0$. We conclude that $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}$ is nonincreasing for all $t\in I$, and therefore $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}\leq\|(w_{1}-w_{2})(\cdot,0)\|_{L^{2}(\Omega)}=0$. Hence $w_{1}\equiv w_{2}$ and the solution (bounded between 0 and $e^{-\lambda\,t}$) is unique.
[…]
2.5 A priori bound
Here we derive a priori bounds with explicit constants.
Theorem 2.2 (a priori bound)
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5). For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10). Then we have the a priori bounds
$\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)},\qquad C^{\boldsymbol{y}}:=\frac{1+a_{\max}^{\boldsymbol{y}}+\lambda^{\boldsymbol{y}}+f_{\max}}{\sqrt{2\,\min(a_{\min}^{\boldsymbol{y}},\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}})}},$
$\|u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\sqrt{\max\big(1+2\,(\lambda^{\boldsymbol{y}})^{2},2\big)}\,e^{\lambda^{\boldsymbol{y}}\,T}\,C^{\boldsymbol{y}}\,\|u_{0}\|_{L^{2}(\Omega)}.$
Recall that the two solutions are related by $u=e^{\lambda\,t}\,w$, with $0\leq u\leq 1$ and $0\leq w\leq e^{-\lambda\,t}$. From PDE II (2.8), we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.18)
Time derivative. Since $0\leq w\leq e^{-\lambda\,t}$, we have $0\leq b\,w^{2}+c\,w\leq(b\,e^{-\lambda\,t}+c)\,w=(\lambda+f)\,w$ and $|(b\,w^{2}+c\,w)v|\leq(\lambda+f)\,w\,|v|$. Thus we obtain from (2.18)
$\Big|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$ $\leq a_{\max}\,\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\lambda+f_{\max})\,\|w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)}\,\|v\|_{H^{1}(\Omega)},$
and therefore
$\|\tfrac{\partial w}{\partial t}(\cdot,t)\|_{H^{1}(\Omega)^{*}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)},$
which yields
$\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}\leq(a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}.$ (2.19)
Gradient. Taking $v=w(\cdot,t)$ in (2.18) and integrating with respect to $t\in I$, we obtain
$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t$
$=\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$
which rearranges into
$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\frac{\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
Since $0\leq w\leq e^{-\lambda\,t}$, we have $(b\,w^{2}+c\,w)\,w\geq c\,w^{2}$. This yields
$\min(a_{\min},c_{\min})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
$\leq a_{\min}\,\int_{I}\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+c_{\min}\,\int_{I}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}^{2}}{2},$
with $c_{\min}:=\lambda-\kappa_{\max}>0$, and therefore
$\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}.$ (2.20)
a priori bound. Combining (2.19) and (2.20), we obtain
$\|w\|_{{\mathcal{X}}}$ $=\sqrt{\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}\leq\|w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{1+a_{\max}+\lambda+f_{\max}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}\,\|w_{0}\|_{L^{2}(\Omega)}.$
Now since $u=e^{\lambda\,t}\,w$, we have
$\|u\|_{{\mathcal{X}}}^{2}$ $=\int_{I}\big(\big\|e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\nabla e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t,$
where we can take out a factor $e^{2\,\lambda\,t}$ from the first two squared norms on the right-hand side, while the third squared norm can be bounded by
$\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}$ $=\big\|\lambda\,e^{\lambda\,t}\,w(\cdot,t)+e^{\lambda\,t}\,\tfrac{\partial}{\partial t}w(\cdot,t)\big\|_{H^{1}(\Omega)^{*}}^{2}$
$\leq 2\,e^{2\lambda\,t}\big(\lambda^{2}\,\|w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}+\|\tfrac{\partial}{\partial t}w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big),$
[…]
3 Parametric regularity
3.1 Uniform random field model
We model the coefficients as affine series expansions in terms of uniform parameters ${\boldsymbol{y}}\in[-\tfrac{1}{2},\tfrac{1}{2}]^{\mathbb{N}}$, see e.g., [21, 29],
$a^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\psi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\psi_{j}({\boldsymbol{x}},t),$ (3.1)
$\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t),$ (3.2)
and both of them are bounded independently of $\boldsymbol{y}$},
$0<a_{\min}:=\|\psi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}$ $<a_{\max}:=\|\psi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<\infty,$
$0<\kappa_{\min}:=\|\xi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}$ $<\kappa_{\max}:=\|\xi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\infty,$
where we define
$\|v\|_{\infty}:=\max_{t\in I}\|v(\cdot,t)\|_{L^{\infty}(\Omega)}.$
To make the two random fields completely independent, we will later define $\psi_{j}\equiv 0$ when $j$ is even and $\xi_{j}\equiv 0$ when $j$ is odd. This can be interpreted as interleaving two series expansions.
Recall it is necessary that $\lambda>\kappa_{\max}\geq 0$, which in this case is independent of $\boldsymbol{y}$}. Then, the coefficients in PDE II are
$b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{\lambda\,t}=\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big)\,e^{\lambda\,t},$ (3.3)
$c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\lambda+f({\boldsymbol{x}},t)-\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)=\lambda+f({\boldsymbol{x}},t)-\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big).$ (3.4)
3.2 Parametric regularity for uniform random fields
We will write derivatives with respect to the parameters $\boldsymbol{y}$} using multiindex notation, where for a multiindex $\boldsymbol{\nu}$$\in{\mathbb{N}}^{\mathbb{N}}$ with $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty$, we define the derivative of order $\boldsymbol{\nu}$} by
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}=\prod_{j\geq 1}\Big(\frac{\partial}{\partial y_{j}}\Big)^{\nu_{j}}.$
We will use the absolute value of the falling factorial as introduced in Chernov and Lê [6, 7]
${[\tfrac{1}{2}]}_{0}:=1,\qquad{[\tfrac{1}{2}]}_{n}:=\big|\tfrac{1}{2}\,(\tfrac{1}{2}-1)\,(\tfrac{1}{2}-2)\cdots(\tfrac{1}{2}-n+1)\big|\quad\mbox{for }n\geq 1,$
with ${[\tfrac{1}{2}]}_{n}\leq n!\leq 2\cdot 2^{n}\,{[\tfrac{1}{2}]}_{n}$, and we will use a key property from [7, Lemma 2.3],
$\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\leq 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|},$ (3.5)
where ${\boldsymbol{m}}<{\boldsymbol{\nu}}$ means that $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ is a product of binomial coefficients.
Theorem 3.1 (parametric regularity)
For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5). For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10). Then for every multiindex $\boldsymbol{\nu}$}, the $\boldsymbol{\nu}$}-th derivative with respect to $\boldsymbol{y}$} of the solutions satisfy the regularity bounds
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)},$
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\widetilde{C}\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\widetilde{C}:=\sqrt{\max(1+2\lambda^{2},2)}\,e^{\lambda\,T}\,\max(C,\tfrac{E}{\rho}),$
with the sequence
$\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty}),$
where $C$ is defined in Theorem 2.2, while $E$ and $\rho$ are defined in (3.20) and (3.21) below, respectively, both depending on $\Phi$ defined in (3.18) below.
For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the mixed derivative $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}$ of (3.1)–(3.4)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}a^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\psi_{j}({\boldsymbol{x}},t)\hskip 14.22636pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.6)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\xi_{j}({\boldsymbol{x}},t)\,e^{\lambda\,t}&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.7)
$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}-\xi_{j}({\boldsymbol{x}},t)\hskip 8.5359pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise}.\end{cases}$ (3.8)
Recall that the weak solution $w\in{\mathcal{X}}$ satisfies (2.12). For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ we take the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the first condition in (2.12) and use the Leibniz product rule to obtain for all $v\in{\mathcal{X}}$,
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $=-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{{\boldsymbol{k}}}a)\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}b)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}c)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$ (3.9)
where the mixed derivative of $w^{2}$ can be expanded further using the Leibniz product rule as
$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{k}}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{k}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}-{\boldsymbol{m}}}w),$
for which the ${\boldsymbol{k}}={\boldsymbol{0}}$ case can be written as
$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)=2\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w+\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w).$
Separating out the ${\boldsymbol{k}}={\boldsymbol{0}}$ terms in (3.2), applying (3.6)–(3.8), and rearranging to have the highest-order mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ on the left-hand side, we obtain
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad+\int_{I}\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.10)
Gradient. Taking the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the second condition in (2.12) and then taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)$, we conclude that $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$. This yields
$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}.$
Taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ in (3.2) then gives
$\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$
Until this point we still have equality in (3.2). For the special cases $|{\boldsymbol{\nu}}|=1$, the sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ is empty and takes the value 0, with no meaning attached to the summand.
We obtain a lower bound on the left-hand side of (3.2) by dropping some nonnegative terms, including the term with integrand $2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\geq 0$ since $0\leq w\leq e^{-\lambda\,t}$, so
LHS $\geq a_{\min}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}^{2}+c_{\min}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
$\geq\min(a_{\min},c_{\min})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2},\qquad c_{\min}:=\lambda-\kappa_{\max}>0.$ (3.12)
For the right-hand side of (3.2), we pull out $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})$ and then apply the Cauchy–Schwarz inequality to the first two terms, as is typically done. The new challenge is to manage the space-time integrals of three factors involving $w$ in the last two terms of (3.2). If we blindly apply the Cauchy–Schwarz inequality multiple times to separate the three factors, then we would raise norms to higher powers than $2$ and the subsequent recursion would fail.
To keep the norms with exponent $2$, we apply the generalized Hölder inequality, first for ${\boldsymbol{x}}\in\Omega$ with $\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1$, and then for $t\in I$ with $\frac{1}{\infty}+\frac{1}{2}+\frac{1}{2}=1$, as follows:
$\int_{I}\int_{\Omega}f\,g\,h\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\int_{I}\|f\|_{L^{2}(\Omega)}\,\|g\|_{L^{4}(\Omega)}\,\|h\|_{L^{4}(\Omega)}\,{\mathrm{d}}t$
$\leq\Big(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\Big)\,\Big(\int_{I}\|g\|_{L^{4}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{L^{4}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
$\leq\theta_{1}^{2}\,\bigg(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\bigg)\,\Big(\int_{I}\|g\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
$=\theta_{1}^{2}\,\|f\|_{C(I;L^{2}(\Omega))}\,\|g\|_{L^{2}(I;H^{1}(\Omega))}\,\|h\|_{L^{2}(I;H^{1}(\Omega))}\leq\theta_{1}^{2}\,\theta_{2}\,\|f\|_{\mathcal{X}}\,\|g\|_{\mathcal{X}}\,\|h\|_{L^{2}(I;H^{1}(\Omega))},$ (3.13)
where we also used the embeddings (2.6) and (2.7). Hence we obtain from (3.2)–(3.2),
$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
$\leq\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\|_{L^{2}(I;L^{2}(\Omega))}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;L^{2}(\Omega))}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))},$
where we used $b\leq\kappa_{\max}\,e^{\lambda\,T}$. The last factor in every term on the right-hand side can be upper bounded by $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$. Canceling out this common factor from both sides and grouping the first two terms on the right-hand side, we arrive at
$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
$\qquad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,=:\,\Theta_{{\boldsymbol{\nu}}}.$ (3.14)
Time derivative. We return to (3.2) but remove the integration over $t\in I$. For a.e. $t\in I$ we have
$\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
$=-\sum_{j\geq 1}\nu_{j}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\sum_{j\geq 1}\nu_{j}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
Using $0\leq w\leq e^{-\lambda\,t}$ to obtain $|2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v+c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v|\leq|\kappa+\lambda+f|\,|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w|\,|v|$, and applying the Cauchy–Schwarz and Hölder inequalities, we obtain
$\Big|\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$
$\leq a_{\max}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\quad+\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}$
$\quad+e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}.$
Using $\|\nabla v\|_{L^{2}(\Omega)}\leq\|v\|_{H^{1}(\Omega)}$ and $\|v\|_{L^{4}(\Omega)}\leq\theta_{1}\,\|v\|_{H^{1}(\Omega)}$, the last factor of every term on the right-hand side can be bounded in terms of $\|v\|_{H^{1}(\Omega)}$. Therefore we conclude that
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{H^{1}(\Omega)^{*}}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{H^{1}(\Omega)}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)},$ (3.15)
where we also used $\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{C(I;L^{2}(\Omega))}\leq\theta_{2}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}$. This step to remove the dependence of one factor on $t$ is crucial before we integrate over $t\in I$.
Next we square both sides of (3.2), integrate over $t\in I$, and then take the square root. We can view the entire right-hand side of (3.2) as a sum of the form $\sum_{k}\alpha_{k}\,g_{k}(t)$, and so we use the estimate
$\bigg(\int_{I}\bigg(\sum_{k}\alpha_{k}\,g_{k}(t)\bigg)^{2}\,{\mathrm{d}}t\bigg)^{1/2}$ $=\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\int_{I}g_{k}(t)\,g_{k^{\prime}}(t)\,{\mathrm{d}}t\bigg)^{1/2}$
$\leq\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg(\int_{I}g_{k^{\prime}}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg)^{1/2}$
$=\sum_{k}\alpha_{k}\,\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}.$
This leads to
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}},$ (3.16)
where $\Theta_{\boldsymbol{\nu}}$ is precisely the right-hand side of (3.2).
Recursion. Using (3.2) and (3.2), we obtain
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $=\sqrt{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}$
$\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}}$
$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\frac{\Theta_{\boldsymbol{\nu}}}{\min(a_{\min},\lambda-\kappa_{\max})}+\Theta_{\boldsymbol{\nu}},$
which yields the recursion
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq\Phi\,\bigg[2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
$\qquad\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\bigg],$ (3.17)
where we abbreviated
$\Phi:=\frac{1+a_{\max}+\kappa_{\max}+\lambda+f_{\max}}{\min(a_{\min},\lambda-\kappa_{\max})}+1.$ (3.18)
Induction. We will prove by induction on $|{\boldsymbol{\nu}}|$ that
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\qquad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}},$ (3.19)
with constants $D$ and $\rho$ to be specified below.
Our base step is $|{\boldsymbol{\nu}}|=1$ (rather than ${\boldsymbol{\nu}}={\boldsymbol{0}}$, as in typical regularity proofs) because our recursion (3.2) involves a sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ which is empty if $|{\boldsymbol{\nu}}|=1$. Recall from Theorem 2.2 that we have the a priori bound $\|w\|_{\mathcal{X}}\leq C\,\|w_{0}\|_{L^{2}(\Omega)}$, where $\|w_{0}\|_{L^{2}(\Omega)}\leq 1$ since $0\leq w_{0}({\boldsymbol{x}})\leq 1$. For ${\boldsymbol{\nu}}={\boldsymbol{e}}_{k}$ for some $k\geq 1$, the recursion (3.2) gives
$\|\partial_{\boldsymbol{y}}^{{\boldsymbol{e}}_{k}}w\|_{\mathcal{X}}$ $\leq\Phi\,\big(2\,\beta_{k}\,\|w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,\|w\|_{\mathcal{X}}^{2}\big)$
$\leq\Phi\,\big(2\,\beta_{k}\,C\,\|w_{0}\|_{L^{2}(\Omega)}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,C^{2}\,\|w_{0}\|_{L^{2}(\Omega)}^{2}\big)\leq\frac{D\,\beta_{k}}{2},$
as required in (3.19), if we define
$D:=2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)\,\|w_{0}\|_{L^{2}(\Omega)}\leq 2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)=:E.$ (3.20)
For the induction step, let $|{\boldsymbol{\nu}}|\geq 2$ and suppose (3.19) is true for all multiindices of order less than $|{\boldsymbol{\nu}}|$. We have from (3.2)
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$
$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}|}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\!\!\!\!\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\!\!\!\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}$
$\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}.$
Using (3.5), this leads to
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\cdot 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,|{\boldsymbol{\nu}}|\cdot\big(2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}\big)$
$\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\big(8+2E\,\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}+16E\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big),$
where assumed $\rho\geq 1$ so that $\rho^{|{\boldsymbol{\nu}}|-3}\leq\rho^{|{\boldsymbol{\nu}}|-2}$, and used $D^{2}\leq DE$ (see (3.20)), together with
$n\,{[\tfrac{1}{2}]}_{n-1}=\frac{n}{|\tfrac{1}{2}-n+1|}\,{[\tfrac{1}{2}]}_{n}\leq\frac{n}{\frac{1}{4}n}\,{[\tfrac{1}{2}]}_{n}=4\,{[\tfrac{1}{2}]}_{n}\quad\mbox{for all }n\geq 2.$
Defining
$\rho:=2\,\Phi\,\big(4+E\,\theta_{1}^{2}\,\theta_{2}\,(\kappa_{\max}+8)\,e^{\lambda\,T}\big)\geq 1,$ (3.21)
we arrive at the required bound (3.19). This completes the induction proof.
We have written the hypothesis (3.19) in a convenient form for induction, where the factor $\rho$ served as a scaling parameter that can be adjusted to suit our needs, as we defined in (3.21). Since $D=E\,\|w_{0}\|_{L^{2}(\Omega)}$ (see (3.20)), this factor $\rho$ can be seen as a rescaling of the sequence $\beta_{j}$ as follows:
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{D}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{E}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}}.$
The result can be extended to cover also the ${\boldsymbol{\nu}}={\boldsymbol{0}}$ case as follows
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}.$
[…]
4 QMC error analysis for uniform random fields
We now fix $\psi_{j}\equiv 0$ for $j$ even and $\xi_{j}\equiv 0$ for $j$ odd to make the two fields independent. We truncate the random fields to $s$ terms and denote the corresponding PDE solution by $u_{s}^{{\boldsymbol{y}}}$ with ${\boldsymbol{y}}=(y_{1},\ldots,y_{s},0,\ldots)$. We are interested in the expected value of the solution $u_{s}$ with respect to the random parameter ${\boldsymbol{y}}\in U_{s}:=[-\frac{1}{2},\frac{1}{2}]^{s}$,
${\mathbb{E}}[u_{s}]:=\int_{U_{s}}u_{s}^{\boldsymbol{y}}\,{\mathrm{d}}{\boldsymbol{y}}.$
We approximate this integral using a randomly shifted lattice rule with generating vector ${\boldsymbol{z}}\in{\mathbb{Z}}^{s}$ and uniformly generated random shift ${\boldsymbol{\Delta}}\in[0,1]^{s}$,
$Q_{N}(u_{s}):=\frac{1}{N}\sum_{i=1}^{N}u^{{\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}}},\quad\mbox{with}\quad{\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}}:=(\tfrac{i}{N}\,{\boldsymbol{z}}+{\boldsymbol{\Delta}})\bmod 1-{\boldsymbol{\tfrac{1}{2}}},$ (4.1)
where the subtraction by ${\boldsymbol{\tfrac{1}{2}}}=(\tfrac{1}{2},\ldots,\tfrac{1}{2})$ translates points from the standard unit cube $0,1]^{s}$ to $U_{s}$. Alternatively, we may wish to compute the expectation of a QoI given by a bounded linear functional $G$ of the PDE solution $u_{s}^{\boldsymbol{y}}$, in which case we approximate ${\mathbb{E}}[G(u_{s})]\approx Q_{N}(G(u_{s}^{\boldsymbol{y}}))$, with the QMC rule applied analogously as in (4.1). For now we work with the exact PDE solution (i.e., no finite element method or time discretization).
Using the parametric regularity bound established in Theorem 3.1, we can apply two general theorems from [21] for Banach-space-valued functions depending on affine uniform parameters $\boldsymbol{y}$}, to obtain a dimension truncation error bound and a QMC error bound. We analyze the QMC error in the first-order weighted unanchored Sobolev space with weight parameters ${\boldsymbol{\gamma}}:=(\gamma_{\mathfrak{u}})_{{\mathfrak{u}}\subset{\mathbb{N}}}$, see e.g., [29].
Theorem 4.1 (dimension truncation error)
Suppose that $\|\psi_{1}\|_{\infty}\geq\|\psi_{3}\|_{\infty}\geq\|\psi_{5}\|_{\infty}\geq\cdots$ and $\|\xi_{2}\|_{\infty}\geq\|\xi_{4}\|_{\infty}\geq\|\xi_{6}\|_{\infty}\geq\cdots$, and $\sum_{j\geq 1}\beta_{j}^{p}=\sum_{j\geq 1}\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})^{p}<\infty$ for some $p\in(0,1)$. Then for all $s\in{\mathbb{N}}$ the dimension truncation error satisfies
$\big\|\,{\mathbb{E}}[u-u_{s}]\,\big\|_{\mathcal{X}}=\Big\|\int_{U}(u^{\boldsymbol{y}}-u^{{\boldsymbol{y}}_{s}})\,{\mathrm{d}}{\boldsymbol{y}}\Big\|_{\mathcal{X}}=\widetilde{C}\,\|u_{0}\|_{L^{2}(\Omega)}\,C^{*}\,s^{-2/p+1},$
where $\widetilde{C}$ is from Theorem 3.1 and $C^{*}>0$ is independent of $s$.
[…]
Theorem 4.2 (QMC error bound)
We have the root-mean-square error bound with respect to the random shift
$\sqrt{{\mathbb{E}}_{\boldsymbol{\Delta}}\,\big[\,\big\|{\mathbb{E}}[u_{s}]-Q_{N}(u_{s})\big\|_{\mathcal{X}}^{2}\big]}\leq C_{s,{\boldsymbol{\gamma}},\lambda}\,[\phi_{\rm tot}(N)]^{-1/(2\lambda)}\quad\mbox{for all }\lambda\in(\tfrac{1}{2},1],$
or for any bounded linear functional $G:{\mathcal{X}}\to{\mathbb{R}}$,
$\sqrt{{\mathbb{E}}_{\boldsymbol{\Delta}}\,\big[\,\big|{\mathbb{E}}[G(u_{s})]-Q_{N}(G(u_{s}))\big|^{2}\big]}\leq\|G\|_{{\mathcal{X}}^{*}}\,C_{s,{\boldsymbol{\gamma}},\lambda}\,[\phi_{\rm tot}(N)]^{-1/(2\lambda)}\quad\mbox{for all }\lambda\in(\tfrac{1}{2},1],$
where $\phi_{\rm tot}(N):=|\{1\leq z\leq N-1:\gcd(z,N)=1\}|$ is the Euler totient function, with $1/\phi_{\rm tot}(N)\leq 2/N$ when $N$ is a prime power, and
$C_{s,{\boldsymbol{\gamma}},\lambda}:=\widetilde{C}\,\bigg(\sum_{\emptyset\neq{\mathfrak{u}}\subseteq\{1:s\}}\gamma_{\mathfrak{u}}^{\lambda}\Big(\frac{2\,\zeta(2\lambda)}{(2\pi^{2})^{\lambda}}\Big)^{|{\mathfrak{u}}|}\bigg)^{1/(2\lambda)}\bigg(\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{{[\tfrac{1}{2}]}_{|{\mathfrak{u}}|}^{2}\,\prod_{j\in{\mathfrak{u}}}(\rho\,\beta_{j})^{2}}{\gamma_{\mathfrak{u}}}\bigg)^{1/2},$
where $\widetilde{C}$ is defined in Theorem 3.1 and $\rho$ is defined in (3.21).
For some $p\in(0,1)$ suppose that $\sum_{j\geq 1}\beta_{j}^{p}=\sum_{j\geq 1}\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})^{p}<\infty$. Then the best convergence rate ${\mathcal{O}}(N^{-\min(1-\delta,\,1/p-1/2)})$ is obtained by the following choices
$\lambda:=\begin{cases}\frac{1}{2-2\delta}\mbox{ for all }\delta\in(0,1)&\mbox{if }p\in[0,\frac{2}{3}],\\ \frac{p}{2-p}&\mbox{if }p\in(\frac{2}{3},1),\end{cases}\quad\gamma_{\mathfrak{u}}:=\bigg({[\tfrac{1}{2}]}_{|{\mathfrak{u}}|}\,\prod_{j\in{\mathfrak{u}}}\frac{\rho\,\beta_{j}}{\sqrt{2\zeta(2\lambda)/(2\pi^{2})^{\lambda}}}\bigg)^{2/(1+\lambda)}.$
and the constant $C_{s,{\boldsymbol{\gamma}},\lambda}$ is bounded independently of $s$.
[…]
5 Numerical experiments
5.1 Experiments with uniform random fields
We consider the spatial domain to be a square with length $L=L_{x_{1}}=L_{x_{2}}=100\,\text{mm}$. The uncertain diffusion and reaction coefficients are defined as in (3.1)–(3.2), taken to be independent of $t$, i.e., stationary in time. The mean fields $a_{0}:=\psi_{0}({\boldsymbol{x}})=0.05$ mm2/day and $\kappa_{0}:=\xi_{0}({\boldsymbol{x}})=0.3$ day-1 are taken to be in realistic ranges based on previous model calibration studies for gliomas [25]. The fluctuations are prescribed by interleaving two sequences as follow:
$\psi_{j}({\boldsymbol{x}})$ $=\begin{cases}a_{0}\,\frac{1}{2}\,k^{-\nu}\sin(\frac{k\,\pi\,x_{1}}{L_{x_{1}}})\,\sin(\frac{k\,\pi\,x_{2}}{L_{x_{2}}})&\mbox{if }j=2k-1.\\ 0&\mbox{if }j=2k,\end{cases}$
$\xi_{j}({\boldsymbol{x}})$ $=\begin{cases}0&\mbox{if }j=2k-1,\\ \kappa_{0}\,\frac{1}{2}\,k^{-\nu}\sin(\frac{k\,\pi\,x_{1}}{L_{x_{1}}})\,\sin(\frac{k\,\pi\,x_{2}}{L_{x_{2}}})&\mbox{if }j=2k,\end{cases}$
for $j\geq 1$ with smoothness parameter $\nu=2$.
The initial condition $u_{0}({\boldsymbol{x}})$ is a Gaussian function in the center of the domain as in [40] and the terminal time is taken to be $T=7$ days. Radiotherapy (2.3) is applied the first five days and chemotherapy (2.4) is administered daily. The radiotherapy dosage is uniformly $z_{\text{rt},k}({\boldsymbol{x}})=2$ Gy and the chemotherapy concentration is assumed to be uniformly $z_{\text{ct},\ell}({\boldsymbol{x}})=1$. The radiosensitivity parameter is $\alpha_{\text{rt}}=0.025\,\text{Gy}^{-1}$ with ratio $\alpha_{\text{rt}}/\beta_{\text{rt}}=10$ Gy following reported values in the literature [40]. The chemotherapy efficacy is $\alpha_{\text{ct}}=0.9$ with clearance rate $\beta_{\text{ct}}=24/1.8~\text{day}^{-1}$. The implicit Euler method is used to discretize (2.1) in time with time step $\Delta t=1/8$ day, such that $\gamma_{\text{rt},\varepsilon}=1/\Delta t$. Linear finite elements are used to discretize the spatial dimension with mesh size $h=1\,\text{mm}$.
We take the quantity of interest to be the total tumor cellularity at the end of the simulation horizon,
$G(u_{s}^{{\boldsymbol{y}}}):=\int_{\Omega}u_{s}^{{\boldsymbol{y}}}({\boldsymbol{x}},T)\,{\mathrm{d}}{\boldsymbol{x}},$
with stochastic dimension $s=256$, that is, there are $128$ terms in the expansion for each field. To obtain an unbiased estimator of ${\mathbb{E}}[G(u_{s}^{\boldsymbol{y}})]$, we average over $R$ i.i.d. uniform random shifts,
$\overline{Q}:=\frac{1}{R}\sum_{r=1}^{R}Q_{N}^{(r)}(G(u_{s})),\qquad Q^{(r)}_{N}[G(u_{s})]:=\frac{1}{N}\sum^{N}_{i=1}G\big(u^{{\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}_{r}}}\big)$,
where, similarly to (4.1), $Q^{(r)}_{N}[G(u_{s})]$ denotes the QMC approximation of the integral ${\mathbb{E}}[G(u_{s})]$ with the $r$th random shift. To estimate the root-mean-square error in Theorem 4.2, we compute the approximation
$\sqrt{\frac{1}{R(R-1)}\sum_{r=1}^{R}\big|\;\overline{Q}-Q_{N}^{(r)}(G(u_{s}))\;\big|^{2}}.$ (5.1)
Here we use $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[4,5,\dots,17]$, for a total of $2^{21}=2,097,152$ PDE evaluations. An off-the-shelf embedded lattice rule with weights tailored to a spectral decay rate of $1/j^{2}$ is utilized to generate the QMC sequence [11]. For comparison, we use $N^{\text{MC}}=2^{17}\cdot R=2^{21}$ quadrature points for MC to match the total number of PDE evaluations. Computations are carried out on the Texas Advanced Computing Center’s Frontera machine [41]. Code to reproduce the experiments is available at: https://github.com/gtpash/qmc-tumor. The results are displayed in Figure 1. We observe that the QMC root-mean-square error converges at a linear rate ${\mathcal{O}}(N^{-1})$ while standard MC converges with the typical $\mathcal{O}(N^{-1/2})$ rate, consistent with the theory.
Figure 1: The approximate root-mean-square error (5.1) of the randomized QMC approximation and the standard error of the MC estimator of the integral ${\mathbb{E}}[G(u_{s})]$ for different numbers of total PDE evaluations.
5.2 Experiments with lognormal random fields
In this section, we consider a more realistic lognormal model for the uncertain coefficients, despite not yet having theoretical results as in the uniform case. The following lognormal model has been successfully employed in Bayesian inversion for gliomas in both murine models [30] and human patients [38],
$a^{\boldsymbol{y}}({\boldsymbol{x}})$ $=\psi_{0}({\boldsymbol{x}})+\exp(Z_{a}({\boldsymbol{x}})),\;Z_{a}({\boldsymbol{x}})=\sum_{j=1}^{\infty}y_{j}\,\psi_{j}({\boldsymbol{x}}),\;\psi_{j}=\begin{cases}\sqrt{\mu_{a,k}}\,\varphi_{a,k}({\boldsymbol{x}})&\mbox{if }j=2k-1,\\ 0&\mbox{if }j=2k,\end{cases}$
$\kappa^{\boldsymbol{y}}({\boldsymbol{x}})$ $=\xi_{0}({\boldsymbol{x}})+\exp(Z_{\kappa}({\boldsymbol{x}})),\;\,Z_{\kappa}({\boldsymbol{x}})=\sum_{j=1}^{\infty}y_{j}\,\xi_{j}({\boldsymbol{x}}),\;\;\xi_{j}=\begin{cases}0&\mbox{if }j=2k-1,\\ \sqrt{\mu_{\kappa,k}}\,\varphi_{\kappa,k}({\boldsymbol{x}})&\mbox{if }j=2k,\end{cases}$
with i.i.d. $y_{j}\sim\mathcal{N}(0,1)$. Here the parameters are stationary in time once more. The infinite sums are the Karhunen–Loève (KL) expansions [32] of the underlying Gaussian random fields $Z_{a}({\boldsymbol{x}})\sim\mathcal{N}(0,\mathcal{C}_{a})$ and $Z_{\kappa}({\boldsymbol{x}})\sim\mathcal{N}(0,\mathcal{C}_{\kappa})$ for the diffusion and proliferation coefficients, with ordered eigenpairs $\{(\mu_{a,k},\varphi_{a,k})\}_{k}$ and $\{(\mu_{\kappa,k},\varphi_{\kappa,k})\}_{k}$, respectively. We again interleave the two independent sequences. The foundational work of [31] established a link between Gaussian random fields with Matérn covariance operators and the solution of a stochastic PDE. Following [42, 47], the covariance operators $\mathcal{C}_{a}$ and $\mathcal{C}_{\kappa}$ are taken to be the inverse of an elliptic operator of the form
$\mathcal{C}:=\mathcal{A}^{-\nu}=(-\gamma\,\Delta+\delta\,I)^{-\nu},$ (5.2)
with smoothness parameter $\nu=2$ and hyperparameters $\gamma>0$ and $\delta>0$ that control the correlation length and pointwise variance [46]. We truncate the expansion and consider the $s$-dimensional subspace spanned by the leading eigenvectors. In particular, we solve a symmetrized (general) eigenvalue problem with a randomized double pass algorithm [47],
$\mathbf{M}\,\mathbf{\Gamma}\,\boldsymbol{\varphi}=\mu\,\mathbf{M}\,\boldsymbol{\varphi},$ (5.3)
where $\mathbf{\Gamma}:=\mathbf{R}^{-1}\mathbf{M}$ is the covariance matrix that arises from the discretization of (5.2), with mass matrix $\mathbf{M}_{ij}=\int_{\Omega}\phi_{i}({\boldsymbol{x}})\phi_{j}({\boldsymbol{x}})\,{\mathrm{d}}{\boldsymbol{x}}$ and bilinear form of the elliptic operator given by $\mathbf{R}_{ij}=\int_{\Omega}\phi_{i}({\boldsymbol{x}})\mathcal{A}^{\nu}\phi_{j}({\boldsymbol{x}})\,{\mathrm{d}}{\boldsymbol{x}}$. Here $\{\phi_{i}\}_{i=1}^{m}$ is the collection of finite element basis functions. The eigenpairs $\{(\mu_{a,k},\varphi_{a,k})\}_{k}$ and $\{(\mu_{\kappa,k},\varphi_{\kappa,k})\}_{k}$ obtained by solving (5.3) may be efficiently computed for generic unstructured domains.
In the numerical experiments, we fix the correlation length to $180$ mm (roughly the largest dimension of a human brain). The pointwise variance is set to $0.2336$ and $0.0682$ for the diffusion and proliferation coefficients, respectively. As in Section 5.1, we take the mean fields to be $a_{0}:=\psi_{0}({\boldsymbol{x}})=0.05$ mm2/day and $\kappa_{0}:=\xi_{0}({\boldsymbol{x}})=0.3$ day-1. The domain is taken to be a two-dimensional slice of a human left-hemisphere [34]. The computational mesh is visualized along with a representative sample in Figure 2. The initial condition $u_{0}({\boldsymbol{x}})$ is once more a Gaussian function centered at the geometric center of the domain with amplitude $0.8$ and width $5.0$ mm. The treatment schedule and radiosensitivity parameters are as in Section 5.1. A time step of $\Delta t=1/8$ day is used for the temporal discretization and the unstructured mesh contains $9,995$ vertices.
We empirically investigate the convergence rate using $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[6,7,\dots 17]$. Each KL expansion is truncated at $512$ terms, for a stochastic dimension of $s=1024$. In the lognormal case, the shifted QMC points are ${\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}}:=\boldsymbol{\Phi}^{-1}\left((\tfrac{i}{N}\,{\boldsymbol{z}}+{\boldsymbol{\Delta}})\bmod 1\right)$, where $\boldsymbol{\Phi}^{-1}(\cdot)$ denotes the inverse cumulative normal distribution applied componentwise to a vector. The same embedded lattice rule as in the uniform case is used to construct the QMC sequence. The empirical spectrum of the covariance operators $\mathcal{C}_{a}$ and $\mathcal{C}_{\kappa}$ and convergence results are shown in Figure 3. We observe that the QMC root-mean-square error converges at a faster rate of $\mathcal{O}(N^{-0.87})$ than the MC rate of $\mathcal{O}(N^{-1/2})$. The square-root of the eigenvalues decay like $k^{-1.17}$ and so we expect the QMC convergence rate to be $1/p-1/2\approx 1.17-1/2=0.67$, if the theory for uniform fields also holds for lognormal fields (as proven for some elliptic linear problems). Hence our results are very encouraging and prompt further analysis.
Figure 2: Left: Visualization of anatomy and slice of the left-hemisphere. Center: The computational mesh used for simulations, with $9,995$ vertices. Right: A representative sample of the diffusion coefficient.
Figure 3: Left: The empirical spectral decay of the two Gaussian random fields compared to the theoretical rate accompanied. Right: The approximate root-mean-square error (5.1) for the QMC approximation and the standard error for the MC estimator of the integral ${\mathbb{E}}[G(u_{s})]$.
6 Conclusions
In this work, we developed a QMC method for a semi-linear parabolic PDE model of tumor growth. The method was applied to the high-dimensional integral arising from estimating the expected value of a smooth QoI. We developed a well-posedness argument for this particular class of problems and derived an a priori bound with an explicit constant. Moreover, in the case of uniform random fields we obtained a parametric regularity bound to provide theoretical error bounds for the QMC approximation. The convergence rate was then numerically verified, in particular, we observe linear convergence. Additionally, we numerically studied application to lognormal random fields and once more observed faster than MC convergence. Future work will focus on extending approximation theory to the lognormal case and application to an optimal control problem. Furthermore, a rigorous analysis of the dimension truncation error as well as the numerical discretization of the PDE should be executed to assess their contributions to the overall discretization error.
Acknowledgments
The authors are grateful for insightful discussions with Albert Cohen, Wolfgang Dahmen, Mike Giles, Maria Pia Gualdani, Pierre L’Ecuyer, Dingcheng Luo, Fabio Nobile, Mark Peletier, William Porteous, and Christian Wieners, as well as other participants at the 2025 Oberwolfach Workshop on Computation and Learning in High Dimensions.

### v2

@@
 The present work extends the theory in the case of non-linear, specifically semi-linear, parabolic equations, providing a roadmap for adapting QMC methods to wider classes of applications.
-Apart from the above work on eigenvalue problems [^3] While eigenvalue problems are also nonlinear, the differential operators [17, 37] are linear., to the best of our knowledge this paper is the first to analyze QMC methods for a non-linear PDE.
+To the best of our knowledge, this paper is the first to analyze QMC methods for a non-linear PDE.
 Well-posedness results are developed for the parametric PDE model in general.
@@
 In particular, it is not monotone and so requires a novel proof, which uses the property that if our initial condition lies in $[0,1]$ then the solution remains in $[0,1]$ (proven in Section 2.5).
-We also makes use of the falling factorial combinatorial bound from [6].
+We also make use of the falling factorial combinatorial bound from [6].
 We stress that obtaining regularity bounds for parabolic (time-dependent) nonlinear problems here is significantly harder than for elliptic problems in [6, 7, 25, 26] due to the need to handle both space and time derivatives and integrals in the norms.
@@
 Combination chemoradiation may be modeled as a linear effect on the state, $f=f({\boldsymbol{x}},t)=f_{\rm rt}({\boldsymbol{x}},t)+f_{\rm ct}({\boldsymbol{x}},t)$, with
-$f_{\rm rt}({\boldsymbol{x}},t)$ $=\gamma_{\text{rt},\varepsilon}\sum_{k}\big[1-\exp(-\alpha_{\text{rt}}\,z_{\text{rt},k}({\boldsymbol{x}})-\beta_{\text{rt}}\,z^{2}_{\text{rt},k}({\boldsymbol{x}}))\big]\,\,\mathbbold{1}(t\in[t_{\text{rt},k},t_{\text{rt},k}+\varepsilon]),$ (2.3)
+$f_{\rm rt}({\boldsymbol{x}},t)$ $=\gamma_{\text{rt},\varepsilon}\sum_{k}\big[1-\exp(-\alpha_{\text{rt}}\,z_{\text{rt},k}({\boldsymbol{x}})-\beta_{\text{rt}}\,z^{2}_{\text{rt},k}({\boldsymbol{x}}))\big]\,\,\mathbbold{1}(t\in[\tau_{\text{rt},k},\tau_{\text{rt},k}+\varepsilon]),$ (2.3)
 $f_{\rm ct}({\boldsymbol{x}},t)$ $=\alpha_{\rm ct}\,\sum_{\ell}z_{\text{ct},\ell}({\boldsymbol{x}})\,\exp(-\beta_{\rm ct}(t-\tau_{{\rm ct},\ell}))\,\mathbbold{1}(t\geq\tau_{\text{ct},\ell}),$ (2.4)
@@
 However, we will restrict the parametric vectors $\boldsymbol{y}$} to some admissible parameter set $U\subset{\mathbb{R}}^{\mathbb{N}}$, such that
-$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}<\infty,\\ 0=f_{\min}\leq f({\boldsymbol{x}},t)\leq f_{\max}<\infty,\end{cases}\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.5)
+$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\,\kappa_{\min}^{\boldsymbol{y}}\leq\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}<\infty,\;\kappa_{\max}^{\boldsymbol{y}}>0,\\ 0=f_{\min}\leq\,f({\boldsymbol{x}},t)\leq f_{\max}<\infty,\end{cases}\qquad\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.5)
 2.2 Functional framework
@@
 The dual norm is defined by
-$\|z\|_{H^{1}(\Omega)^{*}}:=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle z,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}z\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}.$
+$\|z\|_{H^{1}(\Omega)^{*}}:=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle z,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}.$
 It is advantageous to consider functions of time with values in Banach spaces when dealing with time-dependent problems.
@@
 Note that $L^{2}(I;H^{1}(\Omega))^{*}=L^{2}(I;H^{1}(\Omega)^{*})$.
-We follow Tröltzsch [45, §3.4] to analyze our PDE solution in the space $\mathcal{X}$} (denoted by $W(0,T)$ in [45]) which consists of all functions $v\in L^{2}(I;H^{1}(\Omega))$ with (distributional) derivative $\tfrac{\partial}{\partial t}v\in L^{2}(I;H^{1}(\Omega)^{*})$, and we equip the space $\mathcal{X}$} with the (graph) norm
+We follow Tröltzsch [49, §3.4] to analyze our PDE solution in the space $\mathcal{X}$} (denoted by $W(0,T)$ in [49]) which consists of all functions $v\in L^{2}(I;H^{1}(\Omega))$ with (distributional) derivative $\tfrac{\partial}{\partial t}v\in L^{2}(I;H^{1}(\Omega)^{*})$, and we equip the space $\mathcal{X}$} with the (graph) norm
 $\|v\|_{\mathcal{X}}^{2}$ $:=\|v\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}v\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}$
@@
 $\,=\int_{I}\big(\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\nabla v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\tfrac{\partial}{\partial t}v(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t.$
-A fundamental property we will use is that the space $\mathcal{X}$} is embedded in $C(I;L^{2}(\Omega))$, see [45, Theorems 3.10 and 3.11], and therefore we formally write for $v\in{\mathcal{X}}$,
-$\int_{I}\int_{\Omega}\frac{\partial v}{\partial t}v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|v(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
-We will make use of a number of embedding constants which we summarize together here:
+A fundamental property we will use is that the space $\mathcal{X}$} is embedded in $C(I;L^{2}(\Omega))$, see [49, Theorems 3.10 and 3.11].
+We formally write, for $v\in{\mathcal{X}}$ and all $\tau\in I$
+$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}v,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|v(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$ (2.6)
+We will make use of a number of embeddings which we summarize together here:
 $\|v\|_{H^{1}(\Omega)^{*}}$ $\leq\|v\|_{L^{2}(\Omega)}\leq\|v\|_{H^{1}(\Omega)},$
 $\|v\|_{L^{4}(\Omega)}$ $\leq\theta_{1}\,\|v\|_{H^{1}(\Omega)},$ (2.7)
-$\|v\|_{C(I;L^{2}(\Omega))}$ $\leq\theta_{2}\,\|v\|_{\mathcal{X}}.$ (2.7)
-The embedding (2.6) is due to the Sobolev embedding theorem for $H^{1}(\Omega)$, see e.g., [45, Theorem 7.1], which state that $H^{1}(\Omega)$ is embedded in $L^{p}(\Omega)$ when $d=2$ and $1\leq p<\infty$, and when $d\geq 3$ and $1\leq p\leq\frac{2d}{d-2}$.
-Another property we will use is that the space $H^{1}(\Omega)$ is closed under addition, subtraction, and the maximum and minimum operations, see e.g., [16, Lemma 7.6].
+$\|v\|_{C(I;L^{2}(\Omega))}$ $\leq\theta_{2}\,\|v\|_{\mathcal{X}},$ (2.8)
+with $\theta_{1},\theta_{2}>0$.
+The embedding (2.7) is due to the Sobolev embedding theorem for $H^{1}(\Omega)$, see e.g., [49, Theorem 7.1], which states that $H^{1}(\Omega)$ is embedded in $L^{p}(\Omega)$ when $d=2$ and $1\leq p<\infty$, and when $d\geq 3$ and $1\leq p\leq\frac{2d}{d-2}$.
+Another property we will use is that the space $H^{1}(\Omega)$ is closed under addition, subtraction, and the maximum and minimum operations, see e.g., [18, Lemma 7.6].
 This allows us to define continuous piecewise-smooth test functions in $H^{1}(\Omega)$.
 2.3 Reparameterization
-In the case where $f=0$, that is, untreated growth, the PDE model (2.1) is the well-studied Fisher-KPP equation [14, 28], and well-posedness results are readily available, see e.g., [39, §3.10.3].
+In the case where $f=0$, that is, untreated growth, the PDE model (2.1) is the well-studied Fisher-KPP equation [15, 30], and well-posedness results are readily available, see e.g., [42, §3.10.3].
 However, we consider $f\neq 0$, thus standard results are not immediately applicable.
@@
 Consequently, we cannot claim the existence and uniqueness of a solution using known results, such as those in [49].
-Following a standard technique for treating parabolic PDEs involving a linear term with a negative coefficient (see e.g., [13, §7.5 Problem 8]), we use the substitution
+Following a standard technique for treating parabolic PDEs involving a linear term with a negative coefficient (see e.g., [14, §7.5 Problem 8]), we use the substitution
 $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w^{\boldsymbol{y}}({\boldsymbol{x}},t)\,$,
@@
 It follows from (2.5) that
-$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq b^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\kappa_{\max}^{\boldsymbol{y}}\,e^{\lambda^{\boldsymbol{y}}\,T}<\infty,\\ 0<\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}}\leq c^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq\lambda^{\boldsymbol{y}}+f_{\max}-\kappa_{\min}^{\boldsymbol{y}}<\infty,\end{cases}\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.10)
+$\begin{cases}0<a_{\min}^{\boldsymbol{y}}\leq a^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq a_{\max}^{\boldsymbol{y}}<\infty,\\ 0\leq\kappa_{\min}^{\boldsymbol{y}}\leq b_{\min}^{\boldsymbol{y}}\leq b^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq b_{\max}^{\boldsymbol{y}}\leq\kappa_{\max}^{\boldsymbol{y}}\,e^{\lambda^{\boldsymbol{y}}\,T}<\infty,\\ 0<\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}}\leq c_{\min}^{\boldsymbol{y}}\leq c^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq c_{\max}^{\boldsymbol{y}}\leq\lambda^{\boldsymbol{y}}+f_{\max}<\infty,\end{cases}\mbox{for all }{\boldsymbol{x}}\in\Omega,\;t\in I,\;{\boldsymbol{y}}\in U.$ (2.11)
 2.4 Well-posedness
@@
 For fixed ${\boldsymbol{y}}\in U$, a function $u^{\boldsymbol{y}}\in{\mathcal{X}}$ is said to be a weak solution for PDE I (2.1) if
-$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial u}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla u\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(\kappa\,u^{2}+(f-\kappa)\,u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}u(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}u_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.11)
-The first condition was obtained by multiplying the PDE with a test function $v\in{\mathcal{X}}$, integrating over ${\boldsymbol{x}}\in\Omega$ and $t\in I$, applying integration by parts on the diffusion term, and noting that $\int_{\partial\Omega}a\,\nabla u\cdot{\boldsymbol{n}}\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ due to the Neumann boundary condition.
+$\begin{cases}\displaystyle\int_{I}\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\!\int_{I}\!\int_{\Omega}\!\big(a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}u(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}u_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.12)
+The first condition was obtained by testing the PDE with a function $v\in{\mathcal{X}}$, integrating over ${\boldsymbol{x}}\in\Omega$ and $t\in I$, applying integration by parts on the diffusion term, and noting that $\int_{\partial\Omega}a\,\nabla u\cdot{\boldsymbol{n}}\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ due to the Neumann boundary condition.
+The integral $\int_{I}\int_{\Omega}u^{2}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ is well defined because if $u\in{\mathcal{X}}$ then $u^{2}\in L^{2}(I;L^{2}(\Omega))$ via the embedding (2.7).
 Analogously, a function $w^{\boldsymbol{y}}\in{\mathcal{X}}$ is said to be a weak solution for PDE II (2.9) if
-$\begin{cases}\displaystyle\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}w(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}w_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.12)
+$\begin{cases}\displaystyle\int_{I}\big\langle\tfrac{\partial}{\partial t}w,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}\big(a\,\nabla w\cdot\nabla v+(b\,w^{2}+c\,w)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0&\forall\,v\in{\mathcal{X}},\vskip 2.84544pt\\ \displaystyle\int_{\Omega}w(\cdot,0)\,v\,{\mathrm{d}}{\boldsymbol{x}}=\int_{\Omega}w_{0}\,v\,{\mathrm{d}}{\boldsymbol{x}}&\forall\,v\in L^{2}(\Omega).\end{cases}$ (2.13)
 Theorem 2.1 (existence, uniqueness, boundedness)
-For every ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, consider PDE II (2.8) with bounded initial condition $w_{0}\in L^{2}(\Omega)$, $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.9)–(2.10).
-Then, there exists a unique weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE II, which is bounded by $0\leq w^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e.
+For every ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}>0$, consider PDE II (2.9) with bounded initial condition $w_{0}\in L^{2}(\Omega)$, $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.10)–(2.11).
+Then, there exists a unique weak solution $w^{\boldsymbol{y}}\in{\mathcal{X}}$ to PDE II.
+This unique solution is bounded by $0\leq w^{\boldsymbol{y}}({\boldsymbol{x}},t)\leq e^{-\lambda^{\boldsymbol{y}}\,t}$ for a.e.
 ${\boldsymbol{x}}\in\Omega$ and all $t\in I$, and it satisfies the a priori bound $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
@@
   1. 1.
-    For fixed ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}$, we infer from [45, Assumptions 5.1 and 5.2, Lemmas 5.3 and 7.10] that there exists a unique weak solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III with bounded coefficients (2.10), with a priori bound $\|w_{*}^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ for $C^{\boldsymbol{y}}>0$.
+    For fixed ${\boldsymbol{y}}\in U$ and $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}>0$, we infer from [49, Assumptions 5.1 and 5.2, Lemmas 5.3 and 7.10] that there exists a unique weak solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ of PDE III with bounded coefficients (2.11), with a priori bound $\|w_{*}^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ for $C^{\boldsymbol{y}}>0$.
   2. 2.
@@
   4. 4.
-    Finally we prove that this solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ for PDE II is unique, and hence conclude that $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)$ is the unique solution of PDE I, and it satisfies $0\leq u({\boldsymbol{x}},t)\leq 1$ for a.e.
+    Finally we prove that this solution $w_{*}^{\boldsymbol{y}}\in{\mathcal{X}}$ for PDE II is unique.
+In particular, there is no other solution taking values below 0 or above $e^{-\lambda^{\boldsymbol{y}}\,t}$.
+Hence we conclude that $u^{\boldsymbol{y}}({\boldsymbol{x}},t)=e^{\lambda^{\boldsymbol{y}}\,t}\,w_{*}^{\boldsymbol{y}}({\boldsymbol{x}},t)$ is the unique solution of PDE I, and it satisfies $0\leq u({\boldsymbol{x}},t)\leq 1$ for a.e.
 ${\boldsymbol{x}}\in\Omega$ and all $t\in I$.
-The strategy of working with a cutoff function in PDE III and using piecewise test functions (see our proof below) appears already in the literature, sometimes referred to as Stampacchia truncation, see e.g., [4, 10, 39, 45].
+The strategy of working with a cutoff function in PDE III and using continuous piecewise-smooth test functions (see our proof below) appears already in the literature, sometimes referred to as Stampacchia truncation, see e.g., [4, 10, 42, 49].
 However, the problem settings in those references differ from ours, so we include a proof for completeness.
@@
 $d(w)=d({\boldsymbol{x}},t,w):=\begin{cases}0&\mbox{if }w^{\boldsymbol{y}}(x,t)<0,\\ b^{\boldsymbol{y}}({\boldsymbol{x}},t)\,\big(w^{\boldsymbol{y}}({\boldsymbol{x}},t)\big)^{2}+c^{\boldsymbol{y}}({\boldsymbol{x}},t)\,w^{\boldsymbol{y}}(x,t)&\mbox{if }0\leq w^{\boldsymbol{y}}(x,t)\leq e^{-\lambda^{\boldsymbol{y}}\,t},\\ b^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{-2\lambda^{\boldsymbol{y}}\,t}+c^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{-\lambda^{\boldsymbol{y}}\,t}&\mbox{if }w^{\boldsymbol{y}}(x,t)>e^{-\lambda^{\boldsymbol{y}}\,t}.\end{cases}$
-We see that PDE III satisfies [45, Assumption 5.1]: $\Omega$ is a bounded Lipschitz domain; for fixed $w\in{\mathbb{R}}$, the function $d({\boldsymbol{x}},t,w)$ is measurable with respect to ${\boldsymbol{x}}\in\Omega$ and $t\in I$; for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is monotonically increasing (nondecreasing) with respect to $w\in{\mathbb{R}}$.
-We see that PDE III also satisfies [45, Assumption 5.2]: for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is uniformly bounded and Lipschitz in $w\in{\mathbb{R}}$.
-Hence, by [45, Lemma 5.3] and subsequent discussions, there exists a unique weak solution of PDE III in $\mathcal{X}$}.
-Consequently, by [45, Lemma 7.10], this solution satisfies $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
+We see that PDE III satisfies [49, Assumption 5.1]: $\Omega$ is a bounded Lipschitz domain; for fixed $w\in{\mathbb{R}}$, the function $d({\boldsymbol{x}},t,w)$ is measurable with respect to ${\boldsymbol{x}}\in\Omega$ and $t\in I$; for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is monotonically increasing (nondecreasing) with respect to $w\in{\mathbb{R}}$.
+We see that PDE III also satisfies [49, Assumption 5.2]: for all ${\boldsymbol{x}}\in\Omega$ and $t\in I$, the function $d({\boldsymbol{x}},t,w)$ is uniformly bounded and Lipschitz in $w\in{\mathbb{R}}$.
+Hence, by [49, Lemma 5.3] and subsequent discussions, there exists a unique weak solution of PDE III in $\mathcal{X}$}.
+Consequently, by [49, Lemma 7.10], this solution satisfies $\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)}$ with a constant $C^{\boldsymbol{y}}>0$.
 We proceed to prove that this unique weak solution for PDE III is bounded between 0 and $e^{-\lambda\,t}$ if the initial condition is bounded by $0\leq w_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$.
-Integrating (2.13), using integration by parts on the diffusion term and using the zero Neumann boundary condition, we have for a.e.
-$t\in I$ and all $v\in H^{1}(\Omega)$,
-$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}d(w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.15)
-Lower bound. We define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ to be the “unwanted lower part” of $w$,
+Integrating (2.14) with respect to ${\boldsymbol{x}}\in\Omega$, using integration by parts on the diffusion term and using the zero Neumann boundary condition, and then integrating over $t\in[0,\tau]$ for some $\tau\leq T$, we obtain
+$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}w,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\int_{\Omega}\big(a\,\nabla w\cdot\nabla v+d(w)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0\quad\forall\,v\in{\mathcal{X}}.$ (2.16)
+Lower bound. We define $v=w_{\#}\in{\mathcal{X}}$ to be the “unwanted lower part” of $w$,
 $w_{\#}({\boldsymbol{x}},t):=\min(w({\boldsymbol{x}},t),0)=\begin{cases}w({\boldsymbol{x}},t)&\mbox{if }w({\boldsymbol{x}},t)<0,\\ 0&\mbox{otherwise}.\end{cases}$
@@
 If $w({\boldsymbol{x}},t)<0$ then
-$\frac{\partial w}{\partial t}\,w_{\#}=\frac{\partial w_{\#}}{\partial t}\,w_{\#},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=0,$ (2.16)
+$\tfrac{\partial}{\partial t}w=\tfrac{\partial}{\partial t}w_{\#},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=0,$ (2.17)
 where we used Case 1 of (2.15).
 On the other hand, if $w({\boldsymbol{x}},t)\geq 0$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.17) hold trivially.
-Putting (2.16) into (2.15) gives for a.e.
-$t\in I$,
-$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
-We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}$ is a nonincreasing function of $t$.
-Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e.
-${\boldsymbol{x}}\in\Omega$, and in turn this yields
-$w({\boldsymbol{x}},t)\geq 0\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
-Upper bound. Now we define $v=w_{\#}(\cdot,t)\in H^{1}(\Omega)$ on the “unwanted upper part” of $w$,
+Putting (2.17) into (2.16) and using (2.6) gives for all $\tau\in I$,
+$\frac{\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}$
+$=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}w_{\#},w_{\#}\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=-\int_{0}^{\tau}\!\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$
+Thus $\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},\tau)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
+$w({\boldsymbol{x}},\tau)\geq 0\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }\tau\in I.$
+Upper bound. Now we define $v=w_{\#}\in{\mathcal{X}}$ on the “unwanted upper part” of $w$,
 $w_{\#}({\boldsymbol{x}},t):=\max(w({\boldsymbol{x}},t)-e^{-\lambda\,t},0)=\begin{cases}w({\boldsymbol{x}},t)-e^{-\lambda\,t}&\mbox{if }w({\boldsymbol{x}},t)>e^{-\lambda\,t},\\ 0&\mbox{otherwise}.\end{cases}$
@@
 ${\boldsymbol{x}}\in\Omega$, since $w({\boldsymbol{x}},0)=w_{0}({\boldsymbol{x}})\leq 1$, we have $w_{\#}({\boldsymbol{x}},0)=0$ and $\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$.
-If $w({\boldsymbol{x}},t)>e^{-\lambda\,t}$ then $\frac{\partial w_{\#}}{\partial t}=\frac{\partial w}{\partial t}+\lambda\,e^{-\lambda\,t}$, and so
-$\frac{\partial w}{\partial t}\,w_{\#}=\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#},\;\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\;d(w)\,w_{\#}=(\lambda+f)\,e^{-\lambda\,t}\,w_{\#},$ (2.17)
+If $w({\boldsymbol{x}},t)>e^{-\lambda\,t}$ then $\frac{\partial}{\partial t}w_{\#}=\frac{\partial}{\partial t}w+\lambda\,e^{-\lambda\,t}$, and so
+$\tfrac{\partial}{\partial t}w=\tfrac{\partial}{\partial t}w_{\#}-\lambda\,e^{-\lambda\,t},\quad\nabla w\cdot\nabla w_{\#}=|\nabla w_{\#}|^{2},\quad d(w)\,w_{\#}=(\lambda+f)\,e^{-\lambda\,t}\,w_{\#},$ (2.18)
 where we used Case 3 of (2.15).
 On the other hand, if $w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$ then $w_{\#}({\boldsymbol{x}},t)=0$ and the equalities in (2.18) hold trivially.
-Putting (2.17) into (2.15) gives for a.e.
-$t\in I$,
-$\int_{\Omega}\Big(\frac{\partial w_{\#}}{\partial t}-\lambda\,e^{-\lambda\,t}\Big)w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(\lambda+f)\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}},$
-which yields
-$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}=\int_{\Omega}\frac{\partial w_{\#}}{\partial t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}=-\int_{\Omega}a\,|\nabla w_{\#}|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}f\,e^{-\lambda\,t}\,w_{\#}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0.$
-We conclude that $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}^{2}$ is a nonincreasing function of $t$.
-Thus $\|w_{\#}(\cdot,t)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},t)=0$ for a.e.
-${\boldsymbol{x}}\in\Omega$, and in turn this yields
-$w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }t\in I.$
+Putting (2.18) into (2.16) gives for all $\tau\in I$,
+$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}w_{\#}-\lambda\,e^{-\lambda\,t},w_{\#}\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla w_{\#}|^{2}+(\lambda+f)\,e^{-\lambda\,t}\,w_{\#}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
+Using (2.6) then yields
+$\frac{\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}=-\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla w_{\#}|^{2}+f\,e^{-\lambda\,t}\,w_{\#}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$
+Thus $\|w_{\#}(\cdot,\tau)\|_{L^{2}(\Omega)}\leq\|w_{\#}(\cdot,0)\|_{L^{2}(\Omega)}=0$, and hence $w_{\#}({\boldsymbol{x}},\tau)=0$ for a.e. ${\boldsymbol{x}}\in\Omega$, and in turn this yields
+$w({\boldsymbol{x}},\tau)\leq e^{-\lambda\,t}\qquad\mbox{for a.e.\ }{\boldsymbol{x}}\in\Omega\mbox{ and all }\tau\in I.$
 Uniqueness.
@@
 Thus, this solution for PDE III is also a solution for PDE II, demonstrating the existence of a solution for PDE II satisfying $0\leq w({\boldsymbol{x}},t)\leq e^{-\lambda\,t}$.
-Suppose that $w_{1}$ and $w_{2}$ are two solutions for PDE II, both bounded between 0 and $e^{-\lambda\,t}$. Then from (2.8) we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
-$\int_{\Omega}\frac{\partial w_{1}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{1}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{1}^{2}+c\,w_{1})\,v\,{\mathrm{d}}{\boldsymbol{x}},$
-$\int_{\Omega}\frac{\partial w_{2}}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w_{2}\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w_{2}^{2}+c\,w_{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
-Subtracting the two equations, writing $w_{1}^{2}-w_{2}^{2}=(w_{1}+w_{2})\,(w_{1}-w_{2})$, and taking $v=w_{1}(\cdot,t)-w_{2}(\cdot,t)\in H^{1}(\Omega)$, we obtain
-$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}^{2}$
-$=-\int_{\Omega}a\,|\nabla(w_{1}-w_{2})|^{2}\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,(w_{1}+w_{2})+c)\,(w_{1}-w_{2})^{2}\,{\mathrm{d}}{\boldsymbol{x}}\leq 0,$
-where we used $w_{1}({\boldsymbol{x}},t)\geq 0$ and $w_{2}({\boldsymbol{x}},t)\geq 0$.
-We conclude that $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}$ is nonincreasing for all $t\in I$, and therefore $\|(w_{1}-w_{2})(\cdot,t)\|_{L^{2}(\Omega)}\leq\|(w_{1}-w_{2})(\cdot,0)\|_{L^{2}(\Omega)}=0$.
-Hence $w_{1}\equiv w_{2}$ and the solution (bounded between 0 and $e^{-\lambda\,t}$) is unique.
+Suppose that $w_{2}$ is another solution for PDE II, which is not necessarily bounded between 0 and $e^{-\lambda\,t}$.
+Subtracting their respective equations analogous to (2.16), we obtain for all $\tau\in I$,
+$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}(w-w_{2}),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t$
+$+\int_{0}^{\tau}\!\int_{\Omega}\big(a\,\nabla(w-w_{2})\cdot\nabla v+b\,(w^{2}-w_{2}^{2})\,v+c\,(w-w_{2})\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0\quad\forall\,v\in{\mathcal{X}}.$
+Taking $v=w-w_{2}$ and writing $w^{2}-w_{2}^{2}=2\,w\,(w-w_{2})-(w-w_{2})^{2}=2\,w\,v-v^{2}$, we obtain
+$\int_{0}^{\tau}\!\big\langle\tfrac{\partial}{\partial t}v,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla v|^{2}+(2\,b\,w+c)\,v^{2}-b\,v^{3}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
+Since $\|v(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$ and $w\geq 0$, and using (2.6), we conclude that
+$\frac{\|v(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}+a_{\min}\,\|\nabla v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}+c_{\min}\,\|v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}\leq b_{\max}\int_{0}^{\tau}\!\!\int_{\Omega}|v|^{3}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (2.19)
+We now prove that $v\equiv 0$ using a proof by contradiction.
+To facilitate this proof, we introduce the space ${\mathcal{Y}}_{\tau}:=C([0,\tau];L^{2}(\Omega))\cap L^{2}([0,\tau];H^{1}(\Omega))$ for $\tau\in[0,T]$ with the norm
+$\|v\|_{{\mathcal{Y}}_{\tau}}^{2}:=\max_{t\in[0,\tau]}\|v(\cdot,t)\|_{L^{2}(\Omega)}^{2}+\|\nabla v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}+\|v\|_{L^{2}([0,\tau];L^{2}(\Omega))}^{2}.$
+Let $t_{0}:=\max\{\tau\in[0,T]:\|v\|_{{\mathcal{Y}}_{\tau}}=0\}$.
+Then $v(\cdot,t)=w(\cdot,t)-w_{2}(\cdot,t)=0$ for all $t\leq t_{0}$.
+We assume by contradiction that $t_{0}<T$.
+For any $\varepsilon>0$, we take $t_{1}\in(t_{0},T]$ such that $0<\|v\|_{{\mathcal{Y}}_{t_{1}}}<\varepsilon$.
+For any $\tau\leq t_{1}$, the right-hand side of (2.19) can be bounded by $b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{3}$ using (3.2).
+Using this to bound each of the three terms on the left-hand side of (2.19) separately and then combining, we deduce that
+$\|v\|_{{\mathcal{Y}}_{t_{1}}}^{2}\leq\big(2+\tfrac{1}{a_{\min}}+\tfrac{1}{c_{\min}}\big)\,b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{3}<\big(2+\tfrac{1}{a_{\min}}+\tfrac{1}{c_{\min}}\big)\,b_{\max}\,\theta_{1}^{2}\,\theta_{2}\,\,\varepsilon\,\|v\|_{{\mathcal{Y}}_{t_{1}}}^{2},$
 […]
@@
 For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5).
-For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10).
-Then we have the a priori bounds
-$\|w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq C^{\boldsymbol{y}}\,\|w_{0}\|_{L^{2}(\Omega)},\qquad C^{\boldsymbol{y}}:=\frac{1+a_{\max}^{\boldsymbol{y}}+\lambda^{\boldsymbol{y}}+f_{\max}}{\sqrt{2\,\min(a_{\min}^{\boldsymbol{y}},\lambda^{\boldsymbol{y}}-\kappa_{\max}^{\boldsymbol{y}})}},$
-$\|u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\sqrt{\max\big(1+2\,(\lambda^{\boldsymbol{y}})^{2},2\big)}\,e^{\lambda^{\boldsymbol{y}}\,T}\,C^{\boldsymbol{y}}\,\|u_{0}\|_{L^{2}(\Omega)}.$
-Recall that the two solutions are related by $u=e^{\lambda\,t}\,w$, with $0\leq u\leq 1$ and $0\leq w\leq e^{-\lambda\,t}$. From PDE II (2.8), we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
-$\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}$ $=-\int_{\Omega}a\,\nabla w\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}-\int_{\Omega}(b\,w^{2}+c\,w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$ (2.18)
-Time derivative.
-Since $0\leq w\leq e^{-\lambda\,t}$, we have $0\leq b\,w^{2}+c\,w\leq(b\,e^{-\lambda\,t}+c)\,w=(\lambda+f)\,w$ and $|(b\,w^{2}+c\,w)v|\leq(\lambda+f)\,w\,|v|$.
-Thus we obtain from (2.18)
-$\Big|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$ $\leq a_{\max}\,\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\lambda+f_{\max})\,\|w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
-$\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)}\,\|v\|_{H^{1}(\Omega)},$
+Then we have the a priori bound
+$\|u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq C^{\boldsymbol{y}}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\mbox{with}\quad C^{\boldsymbol{y}}:=\frac{(1+a_{\max}^{\boldsymbol{y}}+\kappa_{\max}^{\boldsymbol{y}}+f_{\max})\,e^{\kappa_{\max}^{\boldsymbol{y}}T}}{\sqrt{\min(a_{\min}^{\boldsymbol{y}},\kappa_{\max}^{\boldsymbol{y}})}}.$
+From PDE I (2.1), we have for a.e. $t\in I$ and all $v\in H^{1}(\Omega)$,
+$\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}+\int_{\Omega}\big(a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big)\,{\mathrm{d}}{\boldsymbol{x}}=0.$ (2.20)
+Time derivative. Since $0\leq u\leq 1$, we have $|(\kappa\,u^{2}+(f-\kappa)\,u)\,v|=|\kappa\,u\,(u-1)+f\,u|\,|v|\leq(\kappa+f)\,u\,|v|$. Thus we obtain from (2.20)
+$\big|\big\langle\tfrac{\partial}{\partial t}u,v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\big|$ $\leq a_{\max}\,\|\nabla u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
+$\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{H^{1}(\Omega)}\,\|v\|_{H^{1}(\Omega)},$
 and therefore
-$\|\tfrac{\partial w}{\partial t}(\cdot,t)\|_{H^{1}(\Omega)^{*}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\int_{\Omega}\frac{\partial w}{\partial t}\,v\,{\mathrm{d}}{\boldsymbol{x}}|}{\|v\|_{H^{1}(\Omega)}}\leq(a_{\max}+\lambda+f_{\max})\,\|w(\cdot,t)\|_{H^{1}(\Omega)},$
+$\|\tfrac{\partial u}{\partial t}(\cdot,t)\|_{H^{1}(\Omega)^{*}}=\sup_{0\neq v\in H^{1}(\Omega)}\frac{|\langle\tfrac{\partial}{\partial t}u,v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}|}{\|v\|_{H^{1}(\Omega)}}\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u(\cdot,t)\|_{H^{1}(\Omega)},$
 which yields
-$\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}\leq(a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}.$ (2.19)
-Gradient. Taking $v=w(\cdot,t)$ in (2.18) and integrating with respect to $t\in I$, we obtain
-$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}-\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t$
-$=\int_{I}\int_{\Omega}\frac{\partial w}{\partial t}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=-\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t-\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$
+$\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|u\|_{L^{2}(I;H^{1}(\Omega))}.$ (2.21)
+Gradient. Taking $v=u(\cdot,t)$ in (2.20) and integrating with respect to $t\in[0,\tau]$ for some $\tau\leq T$, and using (2.6), we obtain
+$\frac{\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}-\|u(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{0}^{\tau}\!\!\big(a\,|\nabla u|^{2}+\kappa\,u^{2}+(f-\kappa)\,u\big)\,u\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0,$
 which rearranges into
-$\frac{\|w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla w|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}(b\,w^{2}+c\,w)\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\frac{\|w(\cdot,0)\|_{L^{2}(\Omega)}^{2}}{2}.$
-Since $0\leq w\leq e^{-\lambda\,t}$, we have $(b\,w^{2}+c\,w)\,w\geq c\,w^{2}$. This yields
-$\min(a_{\min},c_{\min})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
-$\leq a_{\min}\,\int_{I}\|\nabla w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+c_{\min}\,\int_{I}\|w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}^{2}}{2},$
-with $c_{\min}:=\lambda-\kappa_{\max}>0$, and therefore
-$\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{\|w_{0}\|_{L^{2}(\Omega)}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}.$ (2.20)
+$\frac{\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}+\!\int_{0}^{\tau}\!\!\int_{\Omega}\big(a\,|\nabla u|^{2}+f\,u^{2}\big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\!\int_{0}^{\tau}\!\!\int_{\Omega}\kappa\,u^{2}\,(1-u)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\frac{\|u_{0}\|_{L^{2}(\Omega)}^{2}}{2}.$ (2.22)
+Dropping the nonnegative second term in (2.22) and using $0\leq u\leq 1$, we obtain the inequality
+$\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}\leq 2\,\kappa_{\max}\int_{0}^{\tau}\|u(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
+The integral form of Grönwall’s inequality (see e.g., [43, Lemma 1.4.1]) then yields
+$\|u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}\leq e^{2\,\kappa_{\max}\,\tau}\|u_{0}\|_{L^{2}(\Omega)}^{2},\quad\mbox{and so}\quad\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}\leq\frac{e^{2\,\kappa_{\max}\,T}-1}{2\,\kappa_{\max}}\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
+Keeping only the $a\,|\nabla u|^{2}$ term on the left-hand side of (2.22) and taking $\tau=T$ lead to
+$a_{\min}\,\|\nabla u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$ $\leq\kappa_{\max}\,\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\frac{\|u_{0}\|_{L^{2}(\Omega)}^{2}}{2}\leq\frac{e^{2\,\kappa_{\max}\,T}}{2}\,\|u_{0}\|_{L^{2}(\Omega)}^{2}.$
+In turn we obtain
+$\|u\|_{L^{2}(I;H^{1}(\Omega))}^{2}=\|u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\|\nabla u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
+$\leq\frac{e^{2\,\kappa_{\max}\,T}-1}{2\,\kappa_{\max}}\|u_{0}\|_{L^{2}(\Omega)}^{2}+\frac{e^{2\,\kappa_{\max}\,T}}{2\,a_{\min}}\,\|u_{0}\|_{L^{2}(\Omega)}^{2}\leq\frac{e^{2\,\kappa_{\max}\,T}}{\min(a_{\min},\kappa_{\max})}\|u_{0}\|_{L^{2}(\Omega)}^{2}.$ (2.23)
 a priori bound. Combining (2.21) and (2.5), we obtain
-$\|w\|_{{\mathcal{X}}}$ $=\sqrt{\|w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}\leq\|w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}w\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
-$\leq(1+a_{\max}+\lambda+f_{\max})\,\|w\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{1+a_{\max}+\lambda+f_{\max}}{\sqrt{2\,\min(a_{\min},\lambda-\kappa_{\max})}}\,\|w_{0}\|_{L^{2}(\Omega)}.$
-Now since $u=e^{\lambda\,t}\,w$, we have
-$\|u\|_{{\mathcal{X}}}^{2}$ $=\int_{I}\big(\big\|e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\nabla e^{\lambda\,t}\,w(\cdot,t)\big\|_{L^{2}(\Omega)}^{2}+\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}\big)\,{\mathrm{d}}t,$
-where we can take out a factor $e^{2\,\lambda\,t}$ from the first two squared norms on the right-hand side, while the third squared norm can be bounded by
-$\big\|\tfrac{\partial}{\partial t}[e^{\lambda\,t}\,w(\cdot,t)]\big\|_{H^{1}(\Omega)^{*}}^{2}$ $=\big\|\lambda\,e^{\lambda\,t}\,w(\cdot,t)+e^{\lambda\,t}\,\tfrac{\partial}{\partial t}w(\cdot,t)\big\|_{H^{1}(\Omega)^{*}}^{2}$
-$\leq 2\,e^{2\lambda\,t}\big(\lambda^{2}\,\|w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}+\|\tfrac{\partial}{\partial t}w(\cdot,t)\|_{H^{1}(\Omega)^{*}}^{2}\big),$
+$\|u\|_{{\mathcal{X}}}=\sqrt{\|u\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}\leq\|u\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}u\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
+$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\|u\|_{L^{2}(I;H^{1}(\Omega))}\leq\frac{(1+a_{\max}+\kappa_{\max}+f_{\max})\,e^{\kappa_{\max}\,T}}{\sqrt{\min(a_{\min},\kappa_{\max})}}\,\|u_{0}\|_{L^{2}(\Omega)},$
 […]
@@
 and both of them are bounded independently of $\boldsymbol{y}$},
-$0<a_{\min}:=\|\psi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}$ $<a_{\max}:=\|\psi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<\infty,$
-$0<\kappa_{\min}:=\|\xi_{0}\|_{\infty}-\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}$ $<\kappa_{\max}:=\|\xi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\infty,$
+$0<a_{\min}:=\min_{t\in I}\inf_{{\boldsymbol{x}}\in\Omega}|\psi_{0}({\boldsymbol{x}},t)|-\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<a_{\max}:=\|\psi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\psi_{j}\|_{\infty}<\infty,$
+$0\leq\kappa_{\min}:=\min_{t\in I}\inf_{{\boldsymbol{x}}\in\Omega}|\xi_{0}({\boldsymbol{x}},t)|-\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\kappa_{\max}:=\|\xi_{0}\|_{\infty}+\frac{1}{2}\sum_{j\geq 1}\|\xi_{j}\|_{\infty}<\infty,$
 where we define
@@
 This can be interpreted as interleaving two series expansions.
-Recall it is necessary that $\lambda>\kappa_{\max}\geq 0$, which in this case is independent of $\boldsymbol{y}$}. Then, the coefficients in PDE II are
-$b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)\,e^{\lambda\,t}=\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big)\,e^{\lambda\,t},$ (3.3)
-$c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $:=\lambda+f({\boldsymbol{x}},t)-\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)=\lambda+f({\boldsymbol{x}},t)-\Big(\xi_{0}({\boldsymbol{x}},t)+\sum_{j\geq 1}y_{j}\,\xi_{j}({\boldsymbol{x}},t)\Big).$ (3.4)
 3.2 Parametric regularity for uniform random fields
-We will write derivatives with respect to the parameters $\boldsymbol{y}$} using multiindex notation, where for a multiindex $\boldsymbol{\nu}$$\in{\mathbb{N}}^{\mathbb{N}}$ with $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty$, we define the derivative of order $\boldsymbol{\nu}$} by
+We will write derivatives with respect to the parameters $\boldsymbol{y}$} using multiindex notation, where for a multiindex ${\boldsymbol{\nu}}\in{\mathbb{N}}^{\mathbb{N}}_{0}$ with $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}<\infty$, we define the derivative of order $\boldsymbol{\nu}$} by
 $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}=\prod_{j\geq 1}\Big(\frac{\partial}{\partial y_{j}}\Big)^{\nu_{j}}.$
@@
 with ${[\tfrac{1}{2}]}_{n}\leq n!\leq 2\cdot 2^{n}\,{[\tfrac{1}{2}]}_{n}$, and we will use a key property from [7, Lemma 2.3],
-$\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\leq 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|},$ (3.5)
-where ${\boldsymbol{m}}<{\boldsymbol{\nu}}$ means that $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ is a product of binomial coefficients.
-Theorem 3.1 (parametric regularity)
-For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5).
-For $\lambda^{\boldsymbol{y}}>\kappa_{\max}^{\boldsymbol{y}}\geq 0$, let $w^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the corresponding solution to PDE II (2.8)–(2.10).
-Then for every multiindex $\boldsymbol{\nu}$}, the $\boldsymbol{\nu}$}-th derivative with respect to $\boldsymbol{y}$} of the solutions satisfy the regularity bounds
-$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}w^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)},$
-$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u^{\boldsymbol{y}}\|_{{\mathcal{X}}}$ $\leq\widetilde{C}\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\widetilde{C}:=\sqrt{\max(1+2\lambda^{2},2)}\,e^{\lambda\,T}\,\max(C,\tfrac{E}{\rho}),$
+$\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\leq 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|},$ (3.3)
+where ${\boldsymbol{m}}<{\boldsymbol{\nu}}$ means that ${\boldsymbol{m}}\neq{\boldsymbol{\nu}}$ and $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ is a product of binomial coefficients.
+Theorem 3.1 (parametric regularity for uniform model)
+For every ${\boldsymbol{y}}\in U$, let $u^{\boldsymbol{y}}\in{\mathcal{X}}$ denote the unique weak solution to PDE I (2.1) with bounded initial condition $u_{0}\in L^{2}(\Omega)$, $0\leq u_{0}({\boldsymbol{x}})\leq 1$ for all ${\boldsymbol{x}}\in\Omega$, and bounded coefficients (2.5) defined by affine expansions (3.1) and (3.2).
+Then for every multiindex $\boldsymbol{\nu}$}, the $\boldsymbol{\nu}$}-th derivative with respect to $\boldsymbol{y}$} of the solution satisfies $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u^{\boldsymbol{y}}\in{\mathcal{X}}$ with the regularity bound
+$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u^{\boldsymbol{y}}\|_{{\mathcal{X}}}\leq\widetilde{C}\,(\rho\,{\boldsymbol{\beta}})^{\boldsymbol{\nu}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)},\quad\widetilde{C}:=\max(C,\tfrac{E}{\rho}),$
 with the sequence
 $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty}),$
-where $C$ is defined in Theorem 2.2, while $E$ and $\rho$ are defined in (3.20) and (3.21) below, respectively, both depending on $\Phi$ defined in (3.18) below.
-For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the mixed derivative $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}$ of (3.1)–(3.4)
-$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}a^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\psi_{j}({\boldsymbol{x}},t)\hskip 14.22636pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.6)
-$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}b^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}\xi_{j}({\boldsymbol{x}},t)\,e^{\lambda\,t}&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise},\end{cases}$ (3.7)
-$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}c^{\boldsymbol{y}}({\boldsymbol{x}},t)$ $=\begin{cases}-\xi_{j}({\boldsymbol{x}},t)\hskip 8.5359pt&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$}\\ 0&\mbox{otherwise}.\end{cases}$ (3.8)
-Recall that the weak solution $w\in{\mathcal{X}}$ satisfies (2.12).
-For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ we take the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the first condition in (2.12) and use the Leibniz product rule to obtain for all $v\in{\mathcal{X}}$,
-$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $=-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{{\boldsymbol{k}}}a)\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}b)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\int_{I}\int_{\Omega}\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}c)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t,$ (3.9)
-where the mixed derivative of $w^{2}$ can be expanded further using the Leibniz product rule as
-$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{k}}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{k}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}-{\boldsymbol{m}}}w),$
+where $C$ is defined in Theorem 2.2 (in the uniform model, $a_{\max}$, $a_{\min}$, $\kappa_{\max}$, $f_{\max}$ are independent of $\boldsymbol{y}$}), while $E$ and $\rho$ are defined in (3.22) and (3.23) below, respectively, both depending on $\Phi$ defined in (3.20) below.
+For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the mixed derivative $\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}$ of (3.1) and (3.2)
+$\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}a^{\boldsymbol{y}}({\boldsymbol{x}},t)=\begin{cases}\psi_{j}({\boldsymbol{x}},t)&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$},\\ 0&\mbox{otherwise},\end{cases}\quad\mbox{and}\quad\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\kappa^{\boldsymbol{y}}({\boldsymbol{x}},t)=\begin{cases}\xi_{j}({\boldsymbol{x}},t)&\mbox{if ${\boldsymbol{\nu}}={\boldsymbol{e}}_{j}$},\\ 0&\mbox{otherwise}.\end{cases}$ (3.4)
+Recall that the weak solution $u\in{\mathcal{X}}$ satisfies (2.12).
+For any multiindex ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ we take the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the first condition in (2.12), but we integrate over $t\in[0,\tau]$ for some $\tau\leq T$ instead of the full interval $I$.
+Using the Leibniz product rule, the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ for the space-time integrand in (2.12) can be written as
+$\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}\big[a\,\nabla u\cdot\nabla v+(\kappa\,u^{2}+(f-\kappa)\,u)\,v\big]$
+$=\sum_{{\boldsymbol{k}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{k}}}\Big[(\partial_{\boldsymbol{y}}^{{\boldsymbol{k}}}a)\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u)\cdot\nabla v+(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}\kappa)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u^{2})\,v+(\partial_{\boldsymbol{y}}^{\boldsymbol{k}}(f-\kappa))\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u)\,v\Big],$ (3.5)
+where the mixed derivative of $u^{2}$ can be expanded further using the Leibniz product rule as
+$\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}}u^{2}$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{k}}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{k}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{k}}-{\boldsymbol{m}}}u),$
 for which the ${\boldsymbol{k}}={\boldsymbol{0}}$ case can be written as
-$(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}}w^{2})$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)=2\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w+\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w).$
-Separating out the ${\boldsymbol{k}}={\boldsymbol{0}}$ terms in (3.2), applying (3.6)–(3.8), and rearranging to have the highest-order mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ on the left-hand side, we obtain
-$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad+\int_{I}\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.10)
-Gradient.
-Taking the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the second condition in (2.12) and then taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)$, we conclude that $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$.
-This yields
-$\int_{I}\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=\int_{I}\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t=\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}.$
-Taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w$ in (3.2) then gives
-$\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,T)\|_{L^{2}(\Omega)}^{2}}{2}+\int_{I}\int_{\Omega}a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)|^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$=-\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t+\sum_{j\geq 1}\nu_{j}\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
-$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{I}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$
-Until this point we still have equality in (3.2).
-For the special cases $|{\boldsymbol{\nu}}|=1$, the sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ is empty and takes the value 0, with no meaning attached to the summand.
-We obtain a lower bound on the left-hand side of (3.2) by dropping some nonnegative terms, including the term with integrand $2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)^{2}\,w\geq 0$ since $0\leq w\leq e^{-\lambda\,t}$, so
-LHS $\geq a_{\min}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}^{2}+c_{\min}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
-$\geq\min(a_{\min},c_{\min})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2},\qquad c_{\min}:=\lambda-\kappa_{\max}>0.$ (3.12)
-For the right-hand side of (3.2), we pull out $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})$ and then apply the Cauchy–Schwarz inequality to the first two terms, as is typically done.
-The new challenge is to manage the space-time integrals of three factors involving $w$ in the last two terms of (3.2).
-If we blindly apply the Cauchy–Schwarz inequality multiple times to separate the three factors, then we would raise norms to higher powers than $2$ and the subsequent recursion would fail.
-To keep the norms with exponent $2$, we apply the generalized Hölder inequality, first for ${\boldsymbol{x}}\in\Omega$ with $\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1$, and then for $t\in I$ with $\frac{1}{\infty}+\frac{1}{2}+\frac{1}{2}=1$, as follows:
+$\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}}u^{2}$ $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)=2\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u+\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u).$
+Separating out the ${\boldsymbol{k}}={\boldsymbol{0}}$ terms in (3.2) and applying (3.4), we obtain for all $v\in{\mathcal{X}}$,
+$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\bigg[a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\cdot\nabla v+2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u\,v+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v$ (3.6)
+$\qquad+\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v+\kappa\,\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,v$
+$\qquad+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\bigg]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0.$
+Membership in $\mathcal{X}$}.
+Taking $\tau=T$ in (3.6) and moving all terms involving derivatives of order strictly less than $\boldsymbol{\nu}$} to the right-hand side, we can see that $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u$ satisfies (the weak form of) a linear parabolic PDE
+$\int_{I}\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t+\int_{I}\int_{\Omega}\big[a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\cdot\nabla v+\widetilde{b}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v\big]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ (3.7)
+$=\int_{I}\langle F_{\boldsymbol{\nu}},v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t,\qquad\forall v\in{\mathcal{X}},$
+where $\widetilde{b}=\widetilde{b}^{\boldsymbol{y}}({\boldsymbol{x}},t)=2\,\kappa\,u+f-\kappa\in L^{2}(I;L^{\infty}(\Omega))$ and $F_{\boldsymbol{\nu}}=F_{\boldsymbol{\nu}}(t)$ is the (time-dependent) functional defined by
+$\langle F_{\boldsymbol{\nu}},v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}:=-\int_{\Omega}\bigg[\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v$ (3.8)
+$+\kappa\,$ $\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,v+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\bigg]\,{\mathrm{d}}{\boldsymbol{x}},$
+which will be shown to be well defined below.
+Since $u_{0}$ is independent of $\boldsymbol{y}$}, differentiating the initial condition in (2.12) gives the initial condition $\int_{\Omega}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0))\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ for all $v\in L^{2}(\Omega)$.
+We now prove by induction that $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$ for all multiindices $\boldsymbol{\nu}$}.
+For the base case ${\boldsymbol{\nu}}={\boldsymbol{0}}$, we have $\partial_{\boldsymbol{y}}^{\boldsymbol{0}}u=u\in{\mathcal{X}}$ by Theorem 2.1.
+Then under the induction hypothesis that $\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\in{\mathcal{X}}$ for all ${\boldsymbol{m}}<{\boldsymbol{\nu}}$, it follows that $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega))^{*}=L^{2}(I;H^{1}(\Omega)^{*})$, which we show below by considering functionals corresponding to each term on the right-hand side of (3.8) separately.
+Let $v\in L^{2}(I;H^{1}(\Omega))$.
+For the first term in (3.8) involving gradients, using $\psi_{j}\in L^{\infty}(\Omega)$ and $\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\in{\mathcal{X}}$ then applying the Cauchy–Schwarz inequality twice gives
+$\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\|\psi_{j}\|_{L^{\infty}(\Omega)}\int_{I}\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}\,{\mathrm{d}}t$
+$\leq\|\psi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{{\mathcal{X}}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
+Hence, the functionals corresponding to the gradient terms in (3.8) are in $L^{2}(I;H^{1}(\Omega)^{*})$. Similarly, for the second term in (3.8) we obtain
+$\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\|\xi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{{\mathcal{X}}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
+For the third and fourth terms in (3.8), we face a new challenge to manage the space-time integrals of three factors involving $u$.
+If we blindly apply the Cauchy–Schwarz inequality multiple times to separate the three factors, then we would raise norms to higher powers than $2$ and the subsequent induction would fail.
+To handle integrals of the form $\int_{I}\int_{\Omega}f\,g\,h\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ where the integrand is a product of three factors and to keep the norms with exponent $2$, we apply the generalized Hölder inequality, first for ${\boldsymbol{x}}\in\Omega$ with $\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1$, and then for $t\in I$ with $\frac{1}{\infty}+\frac{1}{2}+\frac{1}{2}=1$, as follows:
 $\int_{I}\int_{\Omega}f\,g\,h\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\int_{I}\|f\|_{L^{2}(\Omega)}\,\|g\|_{L^{4}(\Omega)}\,\|h\|_{L^{4}(\Omega)}\,{\mathrm{d}}t$
 $\leq\Big(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\Big)\,\Big(\int_{I}\|g\|_{L^{4}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{L^{4}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
-$\leq\theta_{1}^{2}\,\bigg(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\bigg)\,\Big(\int_{I}\|g\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
+$\leq\theta_{1}^{2}\,\Big(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\Big)\,\Big(\int_{I}\|g\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
 $=\theta_{1}^{2}\,\|f\|_{C(I;L^{2}(\Omega))}\,\|g\|_{L^{2}(I;H^{1}(\Omega))}\,\|h\|_{L^{2}(I;H^{1}(\Omega))}\leq\theta_{1}^{2}\,\theta_{2}\,\|f\|_{\mathcal{X}}\,\|g\|_{\mathcal{X}}\,\|h\|_{L^{2}(I;H^{1}(\Omega))},$ (3.9)
-where we also used the embeddings (2.6) and (2.7). Hence we obtain from (3.2)–(3.2),
-$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
-$\leq\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\|_{L^{2}(I;L^{2}(\Omega))}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}$
-$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;L^{2}(\Omega))}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))},$
-where we used $b\leq\kappa_{\max}\,e^{\lambda\,T}$.
-The last factor in every term on the right-hand side can be upper bounded by $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$.
+where we used the embeddings (2.7) and (2.8). Thus we obtain for the third and fourth terms in (3.8),
+$\int_{I}\int_{\Omega}\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))},$
+$\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\theta_{1}^{2}\,\theta_{2}\,\|\xi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
+Seeing as though $F_{\boldsymbol{\nu}}$ is a linear combination of functionals belonging to $L^{2}(I;H^{1}(\Omega)^{*})$, we conclude that $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega)^{*})$.
+Since the bilinear form corresponding to the second term in (3.7) is elliptic and bounded and $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega)^{*})$, it then follows from standard results for linear parabolic PDEs (e.g., [12, 45]) that there exists a unique $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$ satisfying (3.7).
+Gradient.
+Taking the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the second condition in (2.12) and then taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)$, we conclude that $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$.
+Since $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$, by (2.6), this yields
+$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}{\mathrm{d}}t=\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}\geq 0.$
+Taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u$ in (3.6) then gives
+$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}{\mathrm{d}}t+\!\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}\,u+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}+\eta\Big]{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0,$ (3.10)
+where
+$\eta({\boldsymbol{x}},t)$ $:=\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)$ (3.11)
+$\quad+\kappa\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)$
+$\quad+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u).$
+Note that $\eta$ can potentially give both positive and negative values.
+Until this point we still have equality in (3.10)–(3.11).
+For the special cases $|{\boldsymbol{\nu}}|=1$, the sum over ${\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}$ is empty and takes the value 0, with no meaning attached to the summand.
+Recall that $0\leq u\leq 1$.
+Now we drop the nonnegative terms $\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}\,u$ and $f\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}$ in (3.10) and rearrange to obtain the inequality
+$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+\eta\Big]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\kappa_{\max}\int_{0}^{\tau}\!\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t.$
+If we were to return to the derivation of (2.12), but multiply the PDE by $e^{-2\,\kappa_{\max}\,t}$ before integrating over $t\in[0,\tau]$, then the preceding argument would lead to the above inequality but with all integrands multiplied by $e^{-2\,\kappa_{\max}\,t}$, and by the product rule it is equivalent to
+$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\Big(\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,t}\Big)\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+\eta\Big]e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$ (3.12)
+Using $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$ and (2.6), we drop the nonnegative term $a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}$ in (3.12) to obtain
+$\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,\tau}}{2}=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\Big(\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,t}\Big)\,{\mathrm{d}}t\leq\int_{0}^{\tau}\!\!\int_{\Omega}|\eta|\,e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$
+We multiply both sides of the inequality by $2\,e^{2\,\kappa_{\max}\,\tau}$ and integrate over $\tau\in I$ to obtain
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$ $\leq 2\int_{0}^{T}\!\!\int_{0}^{\tau}\!\!\int_{\Omega}|\eta|\,e^{2\,\kappa_{\max}\,(\tau-t)}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\,{\mathrm{d}}\tau=\int_{I}\int_{\Omega}|\eta|\,\frac{e^{2\,\kappa_{\max}\,(T-t)}-1}{\kappa_{\max}}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.13)
+Next we drop the nonnegative first term in (3.12) and take $\tau=T$ to obtain
+$a_{\min}\,e^{-2\,\kappa_{\max}\,T}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}^{2}\leq\int_{I}\int_{\Omega}|\eta|\,e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.14)
+Combining (3.13) and (3.14), we conclude that
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}=\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
+$\leq\int_{I}\int_{\Omega}|\eta|\,\Big(\frac{e^{2\,\kappa_{\max}\,(T-t)}-1}{\kappa_{\max}}+\frac{e^{2\,\kappa_{\max}\,(T-t)}}{a_{\min}}\Big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\frac{2\,e^{2\,\kappa_{\max}\,T}}{\min(a_{\min},\kappa_{\max})}\int_{I}\int_{\Omega}|\eta|\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.15)
+To bound the space-time integral of $|\eta|$ with $\eta$ defined in (3.11), we pull out $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})$ and then apply the Cauchy–Schwarz inequality to the first two terms, as is typically done.
+For the last two terms in (3.11) we again use (3.2) to manage the space-time integrals of three factors.
+We obtain from (3.11), (3.2), and (3.2),
+$\frac{\min(a_{\min},\kappa_{\max})}{2\,e^{2\,\kappa_{\max}\,T}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}\leq\int_{I}\int_{\Omega}|\eta|\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
+$\leq\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\|_{L^{2}(I;L^{2}(\Omega))}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}$
+$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{L^{2}(I;L^{2}(\Omega))}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}$
+$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$
+$\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}.$
+The last factor in every term on the right-hand side can be upper bounded by $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$.
 Canceling out this common factor from both sides and grouping the first two terms on the right-hand side, we arrive at
-$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
-$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
-$\qquad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,=:\,\Theta_{{\boldsymbol{\nu}}}.$ (3.14)
-Time derivative. We return to (3.2) but remove the integration over $t\in I$. For a.e. $t\in I$ we have
-$\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
-$=-\sum_{j\geq 1}\nu_{j}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\sum_{j\geq 1}\nu_{j}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
-$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\$
-$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
-Using $0\leq w\leq e^{-\lambda\,t}$ to obtain $|2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v+c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v|\leq|\kappa+\lambda+f|\,|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w|\,|v|$, and applying the Cauchy–Schwarz and Hölder inequalities, we obtain
-$\Big|\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$
-$\leq a_{\max}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
-$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
-$\quad+\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}$
-$\quad+e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}.$
+$\frac{\min(a_{\min},\kappa_{\max})}{2\,e^{2\,\kappa_{\max}\,T}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$
+$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}$
+$\qquad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,=:\,\Theta_{{\boldsymbol{\nu}}}.$ (3.16)
+Time derivative.
+We return to (3.6) but remove the integration over $t$ and move all terms except the duality pair to the right-hand side.
+Then we apply the triangle, Cauchy–Schwarz and Hölder inequalities, together with $|2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u\,v+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v|\leq(\kappa+f)\,|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u|\,|v|$ since $0\leq u\leq 1$.
+We obtain for a.e.
+$t\in I$,
+$\big|\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\big|$
+$\leq a_{\max}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
+$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
+$\quad+\kappa_{\max}\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}$
+$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}.$
 Using $\|\nabla v\|_{L^{2}(\Omega)}\leq\|v\|_{H^{1}(\Omega)}$ and $\|v\|_{L^{4}(\Omega)}\leq\theta_{1}\,\|v\|_{H^{1}(\Omega)}$, the last factor of every term on the right-hand side can be bounded in terms of $\|v\|_{H^{1}(\Omega)}$.
 Therefore we conclude that
-$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{H^{1}(\Omega)^{*}}$
-$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{H^{1}(\Omega)}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{H^{1}(\Omega)}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)},$ (3.15)
-where we also used $\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{C(I;L^{2}(\Omega))}\leq\theta_{2}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}$.
+$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)(\cdot,t)\|_{H^{1}(\Omega)^{*}}$ $\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{H^{1}(\Omega)}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u(\cdot,t)\|_{H^{1}(\Omega)}$
+$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u(\cdot,t)\|_{H^{1}(\Omega)}$
+$\quad+\theta_{1}^{2}\,\theta_{2}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u(\cdot,t)\|_{H^{1}(\Omega)},$ (3.17)
+where we also used $\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{C(I;L^{2}(\Omega))}\leq\theta_{2}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}$.
 This step to remove the dependence of one factor on $t$ is crucial before we integrate over $t\in I$.
-Next we square both sides of (3.2), integrate over $t\in I$, and then take the square root.
-We can view the entire right-hand side of (3.2) as a sum of the form $\sum_{k}\alpha_{k}\,g_{k}(t)$, and so we use the estimate
-$\bigg(\int_{I}\bigg(\sum_{k}\alpha_{k}\,g_{k}(t)\bigg)^{2}\,{\mathrm{d}}t\bigg)^{1/2}$ $=\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\int_{I}g_{k}(t)\,g_{k^{\prime}}(t)\,{\mathrm{d}}t\bigg)^{1/2}$
-$\leq\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg(\int_{I}g_{k^{\prime}}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg)^{1/2}$
-$=\sum_{k}\alpha_{k}\,\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}.$
-This leads to
-$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
-$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;H^{1}(\Omega))}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
-$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}},$ (3.16)
+Next we take the $L^{2}(I)$ norm of (3.2) and apply the triangle inequality to obtain
+$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$ $\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}},$ (3.18)
 where $\Theta_{\boldsymbol{\nu}}$ is precisely the right-hand side of (3.2).
 Recursion. Using (3.2) and (3.18), we obtain
-$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $=\sqrt{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}$
-$\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
-$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}}$
-$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\frac{\Theta_{\boldsymbol{\nu}}}{\min(a_{\min},\lambda-\kappa_{\max})}+\Theta_{\boldsymbol{\nu}},$
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $=\sqrt{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}$
+$\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
+$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}}$
+$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\frac{2\,e^{2\kappa_{\max}\,T}\,\Theta_{\boldsymbol{\nu}}}{\min(a_{\min},\kappa_{\max})}+\Theta_{\boldsymbol{\nu}},$
 which yields the recursion
-$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq\Phi\,\bigg[2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
-$\qquad\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\bigg],$ (3.17)
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $\leq\Phi\,\bigg[2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}$
+$\qquad\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\bigg],$ (3.19)
 where we abbreviated
-$\Phi:=\frac{1+a_{\max}+\kappa_{\max}+\lambda+f_{\max}}{\min(a_{\min},\lambda-\kappa_{\max})}+1.$ (3.18)
+$\Phi:=\frac{2\,e^{2\,\kappa_{\max}\,T}(1+a_{\max}+\kappa_{\max}+f_{\max})}{\min(a_{\min},\kappa_{\max})}+1.$ (3.20)
 Induction. We will prove by induction on $|{\boldsymbol{\nu}}|$ that
-$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\qquad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}},$ (3.19)
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\qquad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}},$ (3.21)
 with constants $D$ and $\rho$ to be specified below.
-Our base step is $|{\boldsymbol{\nu}}|=1$ (rather than ${\boldsymbol{\nu}}={\boldsymbol{0}}$, as in typical regularity proofs) because our recursion (3.2) involves a sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ which is empty if $|{\boldsymbol{\nu}}|=1$.
-Recall from Theorem 2.2 that we have the a priori bound $\|w\|_{\mathcal{X}}\leq C\,\|w_{0}\|_{L^{2}(\Omega)}$, where $\|w_{0}\|_{L^{2}(\Omega)}\leq 1$ since $0\leq w_{0}({\boldsymbol{x}})\leq 1$.
+Our base step is $|{\boldsymbol{\nu}}|=1$ (rather than ${\boldsymbol{\nu}}={\boldsymbol{0}}$, as in typical proofs of parametric regularity) because our recursion (3.2) involves a sum over ${\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}$ which is empty if $|{\boldsymbol{\nu}}|=1$.
+Recall from Theorem 2.2 that we have the a priori bound $\|u\|_{\mathcal{X}}\leq C\,\|u_{0}\|_{L^{2}(\Omega)}$, where $\|u_{0}\|_{L^{2}(\Omega)}^{2}\leq|\Omega|$ since $0\leq u_{0}({\boldsymbol{x}})\leq 1$.
 For ${\boldsymbol{\nu}}={\boldsymbol{e}}_{k}$ for some $k\geq 1$, the recursion (3.2) gives
-$\|\partial_{\boldsymbol{y}}^{{\boldsymbol{e}}_{k}}w\|_{\mathcal{X}}$ $\leq\Phi\,\big(2\,\beta_{k}\,\|w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,\|w\|_{\mathcal{X}}^{2}\big)$
-$\leq\Phi\,\big(2\,\beta_{k}\,C\,\|w_{0}\|_{L^{2}(\Omega)}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,C^{2}\,\|w_{0}\|_{L^{2}(\Omega)}^{2}\big)\leq\frac{D\,\beta_{k}}{2},$
+$\|\partial_{\boldsymbol{y}}^{{\boldsymbol{e}}_{k}}u\|_{\mathcal{X}}$ $\leq\Phi\,\big(2\,\beta_{k}\,\|u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\beta_{k}\,\|u\|_{\mathcal{X}}^{2}\big)$
+$\leq\Phi\,\big(2\,\beta_{k}\,C\,\|u_{0}\|_{L^{2}(\Omega)}+\theta_{1}^{2}\,\theta_{2}\,\beta_{k}\,C^{2}\,\|u_{0}\|_{L^{2}(\Omega)}\,|\Omega|^{1/2}\big)\leq\frac{D\,\beta_{k}}{2},$
 as required in (3.21), if we define
-$D:=2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)\,\|w_{0}\|_{L^{2}(\Omega)}\leq 2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)=:E.$ (3.20)
+$D:=E\,\|u_{0}\|_{L^{2}(\Omega)},\quad\mbox{with}\quad E:=2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,|\Omega|^{1/2}\big).$ (3.22)
 For the induction step, let $|{\boldsymbol{\nu}}|\geq 2$ and suppose (3.21) is true for all multiindices of order less than $|{\boldsymbol{\nu}}|$. We have from (3.2)
-$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$
+$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$
 $\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}|}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
-$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\!\!\!\!\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\!\!\!\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}$
-$\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
-$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}.$
+$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
+$\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\!\!\!\!\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\!\!\!\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}$
+$\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
+$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}.$
 Using (3.3), this leads to
-$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\cdot 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}$
-$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,|{\boldsymbol{\nu}}|\cdot\big(2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}\big)$
-$\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\big(8+2E\,\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}+16E\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big),$
-where assumed $\rho\geq 1$ so that $\rho^{|{\boldsymbol{\nu}}|-3}\leq\rho^{|{\boldsymbol{\nu}}|-2}$, and used $D^{2}\leq DE$ (see (3.20)), together with
+$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\cdot 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}$
+$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,|{\boldsymbol{\nu}}|\cdot\big(2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}\big)$
+$\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\big(8+2E\,\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,|\Omega|^{1/2}+16E\,\theta_{1}^{2}\,\theta_{2}\,|\Omega|^{1/2}\big),$
+where we assumed $\rho\geq 1$ so that $\rho^{|{\boldsymbol{\nu}}|-3}\leq\rho^{|{\boldsymbol{\nu}}|-2}$, and used $D^{2}\leq DE\,|\Omega|^{1/2}$ (see (3.22)), together with
 $n\,{[\tfrac{1}{2}]}_{n-1}=\frac{n}{|\tfrac{1}{2}-n+1|}\,{[\tfrac{1}{2}]}_{n}\leq\frac{n}{\frac{1}{4}n}\,{[\tfrac{1}{2}]}_{n}=4\,{[\tfrac{1}{2}]}_{n}\quad\mbox{for all }n\geq 2.$
 Defining
-$\rho:=2\,\Phi\,\big(4+E\,\theta_{1}^{2}\,\theta_{2}\,(\kappa_{\max}+8)\,e^{\lambda\,T}\big)\geq 1,$ (3.21)
+$\rho:=2\,\Phi\,\big(4+E\,\theta_{1}^{2}\,\theta_{2}\,(\kappa_{\max}+8)\,|\Omega|^{1/2}\big)\geq 1,$ (3.23)
 we arrive at the required bound (3.21). This completes the induction proof.
 We have written the hypothesis (3.21) in a convenient form for induction, where the factor $\rho$ served as a scaling parameter that can be adjusted to suit our needs, as we defined in (3.23).
-Since $D=E\,\|w_{0}\|_{L^{2}(\Omega)}$ (see (3.20)), this factor $\rho$ can be seen as a rescaling of the sequence $\beta_{j}$ as follows:
-$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{D}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{E}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}}.$
+Since $D=E\,\|u_{0}\|_{L^{2}(\Omega)}$ (see (3.22)), this factor $\rho$ can be seen as a rescaling of the sequence $\beta_{j}$ as follows:
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{D}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{E}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}}.$
 The result can be extended to cover also the ${\boldsymbol{\nu}}={\boldsymbol{0}}$ case as follows
-$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}.$
+$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}.$
 […]
 4 QMC error analysis for uniform random fields
+As noted in the introduction, QMC methods have demonstrated impressive performance in several UQ applications involving PDEs with random coefficients or domains, including linear elliptic PDEs [13, 17, 20, 21, 22, 24, 32], elliptic eigenvalue problems [19, 40] and linear parabolic equations with control [23].
+For a more pedagogical survey of some of the earlier works from the QMC perspective, we refer the interested reader to [31].
 We now fix $\psi_{j}\equiv 0$ for $j$ even and $\xi_{j}\equiv 0$ for $j$ odd to make the two fields independent.
@@
 Alternatively, we may wish to compute the expectation of a QoI given by a bounded linear functional $G$ of the PDE solution $u_{s}^{\boldsymbol{y}}$, in which case we approximate ${\mathbb{E}}[G(u_{s})]\approx Q_{N}(G(u_{s}^{\boldsymbol{y}}))$, with the QMC rule applied analogously as in (4.1).
+A randomly shifted lattice rule is unbiased, i.e., ${\mathbb{E}}_{\boldsymbol{\Delta}}[Q_{N}(G(u_{s}))]={\mathbb{E}}[G(u_{s})]$.
 For now we work with the exact PDE solution (i.e., no finite element method or time discretization).
@@
 $\big\|\,{\mathbb{E}}[u-u_{s}]\,\big\|_{\mathcal{X}}=\Big\|\int_{U}(u^{\boldsymbol{y}}-u^{{\boldsymbol{y}}_{s}})\,{\mathrm{d}}{\boldsymbol{y}}\Big\|_{\mathcal{X}}=\widetilde{C}\,\|u_{0}\|_{L^{2}(\Omega)}\,C^{*}\,s^{-2/p+1},$
-where $\widetilde{C}$ is from Theorem 3.1 and $C^{*}>0$ is independent of $s$.
+where $\widetilde{C}$ is the constant from Theorem 3.1 and $C^{*}>0$ is independent of $s$.
 […]
@@
 The radiosensitivity parameter is $\alpha_{\text{rt}}=0.025\,\text{Gy}^{-1}$ with ratio $\alpha_{\text{rt}}/\beta_{\text{rt}}=10$ Gy following reported values in the literature [44].
-The chemotherapy efficacy is $\alpha_{\text{ct}}=0.9$ with clearance rate $\beta_{\text{ct}}=24/1.8~\text{day}^{-1}$.
-The implicit Euler method is used to discretize (2.1) in time with time step $\Delta t=1/8$ day, such that $\gamma_{\text{rt},\varepsilon}=1/\Delta t$.
-Linear finite elements are used to discretize the spatial dimension with mesh size $h=1\,\text{mm}$.
+The chemotherapy efficacy is taken to be $\alpha_{\text{ct}}=0.9$ with clearance rate $\beta_{\text{ct}}=24\,\ln(2)/1.8~\text{day}^{-1}$ for a chemotherapy based on Temozolomide, which has a half-life of 1.8 hours [16].
+The implicit Euler method is used to discretize (2.1) in time with time step $\Delta t=1/8$ day, such that $\gamma_{\text{rt},\varepsilon}=1/\Delta t$, where $\varepsilon=\Delta t$ matches the resolution of the time-discretization.
+From (2.3), for each $k$, we compute the time-integrated effect of one radiotherapy dose
+$\int f_{\text{rt},k}\,dt=1-\exp(-\alpha_{\text{rt}}z_{\text{rt},k}-\beta_{\text{rt}}z^{2}_{\text{rt},k})\approx 0.06,$
+hence each dose of radiotherapy reduced $u$ by about $6\%$. Similarly, for each $\ell$,
+$\int f_{\text{ct},\ell}\,dt=\alpha_{\text{ct}}z_{\text{ct},\ell}\int_{\tau_{\text{ct},\ell}}^{\infty}\exp(-\beta_{\text{ct}}(t-\tau_{\text{ct},\ell}))=\frac{\alpha_{\text{ct}}\,z_{\text{ct},\ell}}{\beta_{\text{ct}}}\approx 0.10,$
+and so each dose of chemotherapy reduces $u$ by about $10\%$, with a slight time delay. Linear finite elements are used to discretize the spatial dimension with mesh size $h=1\,\text{mm}$.
 We take the quantity of interest to be the total tumor cellularity at the end of the simulation horizon,
@@
 with stochastic dimension $s=256$, that is, there are $128$ terms in the expansion for each field.
-To obtain an unbiased estimator of ${\mathbb{E}}[G(u_{s}^{\boldsymbol{y}})]$, we average over $R$ i.i.d.
+To obtain an unbiased estimator of ${\mathbb{E}}[G(u_{s})]=\int_{U_{s}}G(u_{s}^{\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}$, we average over $R$ i.i.d.
 uniform random shifts,
 $\overline{Q}:=\frac{1}{R}\sum_{r=1}^{R}Q_{N}^{(r)}(G(u_{s})),\qquad Q^{(r)}_{N}[G(u_{s})]:=\frac{1}{N}\sum^{N}_{i=1}G\big(u^{{\boldsymbol{y}}^{(i)}_{{\boldsymbol{\Delta}}_{r}}}\big)$,
-where, similarly to (4.1), $Q^{(r)}_{N}[G(u_{s})]$ denotes the QMC approximation of the integral ${\mathbb{E}}[G(u_{s})]$ with the $r$th random shift.
+where, similarly to (4.1), $Q^{(r)}_{N}[G(u_{s})]$ denotes the unbiased QMC approximation of the integral ${\mathbb{E}}[G(u_{s})]$ with the $r$th random shift.
+Therefore the average $\overline{Q}$ is also an unbiased estimator of ${\mathbb{E}}[G(u_{s})]$, and its variance is $1/R$ times the variance for any individual $Q^{(r)}_{N}[G(u_{s})]$.
 To estimate the root-mean-square error in Theorem 4.2, we compute the approximation
@@
 For comparison, we use $N^{\text{MC}}=2^{17}\cdot R=2^{21}$ quadrature points for MC to match the total number of PDE evaluations.
-Computations are carried out on the Texas Advanced Computing Center’s Frontera machine [41].
+Computations are carried out on the Texas Advanced Computing Center’s Lonestar6 cluster across two compute nodes with 256 total parallel MPI processes, each evaluating one PDE.
+The MC reference calculation with $2^{21}$ evaluations took 1,141 minutes ($\approx 19$ hours), while the QMC shift replicates with $2^{17}$ evaluations averaged 146 minutes ($\approx 2.4$ hours).
 Code to reproduce the experiments is available at: https://github.com/gtpash/qmc-tumor.
@@
 In this section, we consider a more realistic lognormal model for the uncertain coefficients, despite not yet having theoretical results as in the uniform case.
+There is QMC theory to handle the lognormal case, see e.g., [20, 21], but it requires a new parametric regularity bound for the lognormal model of this problem which is anticipated to be harder and so is left for future work.
 The following lognormal model has been successfully employed in Bayesian inversion for gliomas in both murine models [33] and human patients [41],
@@
 The same embedded lattice rule as in the uniform case is used to construct the QMC sequence.
+Computations are carried out in the same fashion as in Section 5.1.
+The reference calculation took 1,417 minutes (23.6 hours) and the QMC evaluations averaged 165 minutes (2.8 hours).
+We attribute the slight increase in computational cost to the larger stochastic dimension which resolves finer features as well as the non-uniform domain.
 The empirical spectrum of the covariance operators $\mathcal{C}_{a}$ and $\mathcal{C}_{\kappa}$ and convergence results are shown in Figure 3.
-We observe that the QMC root-mean-square error converges at a faster rate of $\mathcal{O}(N^{-0.87})$ than the MC rate of $\mathcal{O}(N^{-1/2})$.
+We observe that the QMC root-mean-square error converges at a faster rate of $\mathcal{O}(N^{-0.86})$ than the MC rate of $\mathcal{O}(N^{-1/2})$.
 The square-root of the eigenvalues decay like $k^{-1.17}$ and so we expect the QMC convergence rate to be $1/p-1/2\approx 1.17-1/2=0.67$, if the theory for uniform fields also holds for lognormal fields (as proven for some elliptic linear problems).
@@
 Acknowledgments
-The authors are grateful for insightful discussions with Albert Cohen, Wolfgang Dahmen, Mike Giles, Maria Pia Gualdani, Pierre L’Ecuyer, Dingcheng Luo, Fabio Nobile, Mark Peletier, William Porteous, and Christian Wieners, as well as other participants at the 2025 Oberwolfach Workshop on Computation and Learning in High Dimensions.
+The authors give heartfelt thanks to both referees for their detailed suggestions to improve the paper, especially to one referee for explaining how to prove that the unique weak solution is always nonnegative.
+We are also grateful for insightful discussions with Albert Cohen, Wolfgang Dahmen, Mike Giles, Maria Pia Gualdani, Pierre L’Ecuyer, Dingcheng Luo, Fabio Nobile, Mark Peletier, William Porteous, and Christian Wieners, as well as other participants at the 2025 Oberwolfach Workshop on Computation and Learning in High Dimensions.